### v1

# On Secret-Message Transmission by Echoing Encrypted Probes

Abstract
A scheme for secure communications, called “Secret-message Transmission by Echoing Encrypted Probes (STEEP)”, is revisited. STEEP is a round-trip scheme with a probing phase from one user to another and an echoing phase in the reverse direction. STEEP is broadly applicable to yield a positive secrecy rate in bits per channel use even if the receive channels at eavesdropper (Eve) are stronger than those between legitimate users in both directions. Differing from prior cases of STEEP, this paper presents STEEP in each of the following settings: using Gaussian channel probing (GCP) and Gaussian linear encryption (GLE) over MIMO channels; using phase-shift-keying (PSK) channel probing and PSK nonlinear encryption over SISO channels; and using GCP and GLE over multiple access channels. In each setting, Eve is assumed to have any number of antennas, and STEEP is shown to yield a positive secrecy rate subject to a sufficiently large power in the echoing phase relative to the power in the probing phase, as long as Eve’s receive channel in the probing phase is not noiseless. Also shown is that the GCP-GLE based STEEP for two users, subject to a form of large powers from the users, has its secrecy rate approaching the secret-key capacity based on GCP.
Index Terms:
Secure communications, information security, secret-key generation, secret-message transmission
I Introduction
Secret-message transmission from one node to another subject to eavesdropping has been a long-standing problem for secure communications, which is encountered widely in modern networks. The information-theoretical study of this problem, nowadays known as physical layer security, has a long history since Shannon’s work [1] in 1940’s. Comprehensive reviews of this subject are available in [2], [3] and [4] among others. Many achievements of great importance have been made by researchers in this field, which are centered around wiretap channel (WTC) and secret key generation (SKG). Yet, to the author’s knowledge, few among the numerous works on WTC developed since [5] and [6] in 1970’s could tell us how to produce a positive secrecy rate between Alice and Bob when the channel between them is half-duplex and always weaker than the receive channel at an eavesdropper (Eve). And few among the numerous works on SKG developed since [7] and [8] in 1990’s could tell us how to connect their developments to a WTC scheme in a broadly beneficial way. There appears a non-negligible disconnect between the numerous works on WTC and those on SKG.
A notable exception is however the work in [22] where a two-way protocol for a binary symmetric channel system is proposed to achieve a positive secrecy rate even if Eve’s channel is stronger than users’. In fact, there is a general principle that predates and underpins this protocol, which consists of two integral steps:
First, if Alice transmits independent realizations of a random integer $X$ over a (memoryless) WTC system, and Bob and Eve receive the corresponding realizations of the random integers (binary or not) $Y$ and $Z$, then it is known [2] that the secret-key capacity $C_{key}$ in bits per realization of $\{X,Y,Z\}$ achievable by Alice and Bob via public communications satisfies
$I(X;Y)-I(Y;Z)\leq C_{key}\leq I(X;Y|Z),$ (1)
where $I(X;Y|Z)$ (for example) denotes the mutual information between $X$ and $Y$ conditional on $Z$. This is also known as Maurer’s bounds [7]. In some cases, the upper and lower bounds coincide, i.e., $C_{key}=I(X;Y)-I(Y;Z)=I(X;Y|Z)$ (e.g., see [11] for a Gaussian case), which is generally positive regardless of the WTC secrecy capacity $[I(X;Y)-I(X;Z)]^{+}$ from Alice to Bob. Here $x^{+}\doteq\max(x,0)$.
Second, given the random integers $X$, $Y$ and $Z$ at Alice, Bob and Eve respectively, an encryption lemma (see section 4.2.1 in [2] or section 22.4.3 in [28]) says that Bob can choose a uniform random integer $S$ and transmit $S\oplus Y$ (a modulo sum of $S$ and $Y$) via a public channel so that the secrecy rate of the effective WTC system from Bob to Alice equals $I(X;Y)-I(Y;Z)$.
The above two-step principle is also a foundation for a scheme called “Secret-message Transmission by Echoing Encrypted Probes (STEEP)” [10]. However, differing from [7], [2], [28] and [22], STEEP as shown in this paper allows the following extensions: $X$, $Y$ and $Z$ are allowed to be real, complex, vectors and/or matrices; the modulo sum $\oplus$ is allowed to be replaced by other suitable operations (examples will be shown); and the public channel from Bob to Alice and Eve may be replaced by any channels at the physical (or an upper) layer. Not necessarily all optimal in information theory, these extensions allow secure communications in a wider range of settings to be conducted rather simply with a guaranteed positive secrecy rate.
Figure 1: STEEP is a round-trip scheme with embedded secret message on returned (estimated) probes.
As illustrated in Fig. 1, there are two collaborative phases in STEEP. In phase 1 (or probing phase), random symbols or probes are transmitted from Alice to Bob. These probes arrive at Bob after a transformation by the channel response, which could result in some “effective” probes that can be estimated consistently (but not necessarily perfectly) by Bob. The exact definition of “effective probe” may vary, depending on how STEEP is implemented. In phase 2 (or echoing phase) of STEEP, Bob’s estimates of the effective probes are encrypted or combined with secret message symbols before they are transmitted (“echoed” back) to Alice. These collaborative two-phase operations result in an effective WTC system from Bob to Alice and Eve, which is almost surely in favor of the users subject to a sufficient power from Bob.
Evolved from a scheme (called iSAT) shown in [9], STEEP is a collaborative round-trip scheme between half-duplex nodes, which has a broad applicability and differs from many two-way full-duplex schemes in the literature such as [27], [17], [15] and [16]. A latest work in [25] also assumes two-way full-duplex to explore the fundamental limits on the secrecy rate region between two users with little consideration of complexity and practicality. The scheme in [22] however can be seen as a special case of STEEP subject to binary symmetric channels.
The goal of this paper is to present STEEP in its latest forms. The primary contributions include novel insights into STEEP in three different settings. The first setting (or G-STEEP) uses Gaussian channel probing (GCP) and Gaussian linear encryption (GLE) over MIMO channels between two users, for which an achievable secrecy rate $C_{s,G}$ is derived and analyzed. In particular, $C_{s,G}$ is shown to converge to the secret-key capacity $C_{key}$ based on GCP over MIMO channels if the echoing power in G-STEEP dominates the probing power and both become large. The second setting (or P-STEEP) uses phase-shift-key (PSK) channel probing and PSK nonlinear encryption between two users, for which an achievable secrecy rate $C_{s,P}$ is also presented. The third setting (or M-STEEP) uses GCP and GLE over multiple access channels between an access point (AP) and multiple users all of whom apply the same probes from the AP. An achievable secrecy rate $C_{s,1}$ of M-STEEP from an arbitrary user to AP is shown to be a function that decreases gradually with some robustness (instead of abruptly) as the number $M$ of users increases. In each setting, the achievable secrecy rate of STEEP is shown to be positive subject to a sufficiently large power in the echoing phase. Note that “capacity” and “achievable rate” are treated interchangeably in this paper since each stated achievable rate is the maximum possible under stated conditions.
The paper is organized as follows. The physical-layer channel models of interest in this paper are described in section II. G-STEEP, P-STEEP and M-STEEP are presented and analyzed respectively in sections III, IV and V. Much of the technical proofs is relegated to the appendix. The paper is ended with additional comments and conclusion.
II Channel Models
Assume that the numbers of antennas on Alice, Bob and Eve are respectively $n_{A}$, $n_{B}$ and $n_{E}$. In the case of wireline communications, each antenna here corresponds to a transceiver.
When Alice transmits (within a coherence time $\mathcal{T}_{1}$) a sequence of random vectors $\sqrt{\frac{p_{A}}{n_{A}}}\mathbf{x}_{A}(k)\in\mathcal{C}^{n_{A}\times 1}$ of power $p_{A}$, we assume that Bob and Eve receive respectively
$\mathbf{y}_{B}(k)=\sqrt{p_{A}/n_{A}}\mathbf{H}_{BA}\mathbf{x}_{A}(k)+\mathbf{w}_{B}(k),$ (2)
$\mathbf{y}_{EA}(k)=\sqrt{p_{A}/n_{A}}\mathbf{H}_{EA}\mathbf{x}_{A}(k)+\mathbf{w}_{EA}(k).$ (3)
where all noises are normalized circular complex Gaussian noises, i.e., $\mathbf{w}_{B}(k)$ is $\mathcal{CN}(0,\mathbf{I}_{n_{B}})$ and $\mathbf{w}_{EA}(k)$ is $\mathcal{CN}(0,\mathbf{I}_{n_{E}})$. For notational simplicity, we will also use the scaled versions of $\mathbf{H}_{BA}$ and $\mathbf{H}_{EA}$, i.e., $\mathbf{H}_{BA}^{\prime}\doteq\sqrt{p_{A}/n_{A}}\mathbf{H}_{BA}\in\mathcal{C}^{n_{B}\times n_{A}}$ and $\mathbf{H}_{EA}^{\prime}\doteq\sqrt{p_{A}/n_{A}}\mathbf{H}_{EA}\in\mathcal{C}^{n_{E}\times n_{A}}$. Here $k$ is the sampling index.
Similarly, when Bob transmits (within a coherence time $\mathcal{T}_{2}$) a sequence of random vectors $\sqrt{\frac{p_{B}}{n_{B}}}\mathbf{x}_{B}(k)\in\mathcal{C}^{n_{B}\times 1}$ of power $p_{B}$, we assume that Alice and Eve receive respectively
$\mathbf{y}_{A}(k)=\sqrt{p_{B}/n_{B}}\mathbf{H}_{AB}\mathbf{x}_{B}(k)+\mathbf{w}_{A}(k),$ (4)
$\mathbf{y}_{EB}(k)=\sqrt{p_{B}/n_{B}}\mathbf{H}_{EB}\mathbf{x}_{B}(k)+\mathbf{w}_{EB}(k),$ (5)
where the normalized noises $\mathbf{w}_{A}(k)$ and $\mathbf{w}_{EB}(k)$ are $\mathcal{CN}(\mathbf{0},\mathbf{I}_{n_{A}})$ and $\mathcal{CN}(\mathbf{0},\mathbf{I}_{n_{E}})$. We will also write $\mathbf{H}_{AB}^{\prime}=\sqrt{p_{B}/n_{B}}\mathbf{H}_{AB}\in\mathcal{C}^{n_{A}\times n_{B}}$ and $\mathbf{H}_{EB}^{\prime}=\sqrt{p_{B}/n_{B}}\mathbf{H}_{EB}\in\mathcal{C}^{n_{E}\times n_{B}}$.
Alice and Bob are half-duplex. Namely, $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ do not overlap. But $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ may or may not belong to a common coherence period.
Each receive channel state information (CSI) is assumed to be known to the corresponding receiver. If there is any CSI feedback between Alice and Bob, this CSI is also assumed to be known to Eve. In fact, all CSI in this paper is treated as known to Eve.
Also assume that all signals and noises in each transmission direction (i.e., from Alice to Bob, or from Bob to Alice) are temporally independent. So, for simpler notations, we will also drop the sampling (or slot) index “$k$”. In this case, one should view the channel matrices as constant but the transmitted signals (and the noises) as random. The results on secrecy rates will be based on a large number of slots in each of probing and echoing phases. In the case of temporally coded transmissions, the assumption of “temporal independence” could typically serve as an approximation.
III STEEP with Gaussian Channel Probing and Gaussian Linear Encryption (G-STEEP)
In this section, G-STEEP is presented, and an achievable secrecy rate $C_{s,G}$ of G-STEEP is also derived and discussed. Properties of $C_{s,G}$ subject to large powers are highlighted.
Given $n_{A}\geq n_{B}$, the probing phase should be from Alice to Bob in order to have the largest $C_{s,G}$. This is because the degree of freedom (DoF) of the secret-key capacity $C_{key}$ based on channel probing from a node with more antennas is larger than that from a node with less antennas. See [9] and [11]. Such a connection between $C_{s,G}$ and $C_{key}$ will also be shown.
III-A Phase 1 of G-STEEP
In phase 1, Alice applies Gaussian probing, i.e., she transmits a realization of the random probing vector $\sqrt{\frac{p_{A}}{n_{A}}}\mathbf{x}_{A}$ in each probing slot where $\mathbf{x}_{A}$ is assumed to be $\mathcal{CN}(0,\mathbf{I}_{n_{A}})$. The corresponding signal received by Bob is $\mathbf{y}_{B}$ in (4) (with “$(k)$” dropped). Let the (thin) SVD of $\mathbf{H}_{BA}$ be $\mathbf{U}_{BA}\boldsymbol{\Pi}_{BA}\mathbf{V}_{BA}^{H}=\sum_{i=1}^{n_{B}}\pi_{BA,i}\mathbf{u}_{BA,i}\mathbf{v}_{BA,i}^{H}$ where $\mathbf{U}_{BA}=[\mathbf{u}_{BA,1},\cdots,\mathbf{u}_{BA,n_{B}}]\in\mathcal{C}^{n_{B}\times n_{B}}$ is unitary and $\mathbf{V}_{BA}=[\mathbf{v}_{BA,1},\cdots,\mathbf{v}_{BA,n_{B}}]\in\mathcal{C}^{n_{A}\times n_{B}}$ is column-wise orthonormal. Then we can write
$\mathbf{y}_{B}=\mathbf{U}_{BA}\boldsymbol{\Pi}_{BA}^{\prime}\mathbf{p}+\mathbf{w}_{B}$ (6)
with $\boldsymbol{\Pi}_{BA}^{\prime}=\sqrt{p_{A}/n_{A}}\boldsymbol{\Pi}_{BA}$. Here $\mathbf{p}\doteq\mathbf{V}_{BA}^{H}\mathbf{x}_{A}$ which is here by definition the effective probing vector at Bob. Clearly, $\mathbf{p}$ is $\mathcal{CN}(0,\mathbf{I}_{n_{B}})$ given $\mathbf{x}_{A}$ being $\mathcal{CN}(0,\mathbf{I}_{n_{A}})$.
Given the Gaussian signal and noise model, the MMSE (minimum-mean-squared-error) estimate $\mathbf{\hat{p}}$ of $\mathbf{p}$ by Bob is linear and given by
$\mathbf{\hat{p}}=\mathbb{E}\{\mathbf{p}\mathbf{y}_{B}^{H}\}(\mathbb{E}\{\mathbf{\mathbf{y}_{B}}\mathbf{y}_{B}^{H}\})^{-1}\mathbf{y}_{B}$
$=\boldsymbol{\Pi}_{BA}^{\prime}\mathbf{U}_{BA}^{H}\left(\mathbf{U}_{BA}\boldsymbol{\Pi}_{BA}^{\prime 2}\mathbf{U}_{BA}^{H}+\mathbf{I}_{n_{B}}\right)^{-1}\mathbf{y}_{B}$
$=\boldsymbol{\Pi}_{BA}^{\prime}\left(\boldsymbol{\Pi}_{BA}^{\prime 2}+\mathbf{I}_{n_{B}}\right)^{-1}\mathbf{U}_{BA}^{H}\mathbf{y}_{B},$ (7)
and $\mathbf{R}_{\mathbf{\hat{p}}}\doteq\mathbb{E}\{\mathbf{\hat{p}}\mathbf{\hat{p}}^{H}\}=\boldsymbol{\Pi}_{BA}^{\prime 2}(\boldsymbol{\Pi}_{BA}^{\prime 2}+\mathbf{I}_{n_{B}})^{-1}$. The operator $\mathbb{E}$ denotes the expectation. The MSE matrix of $\mathbf{\hat{p}}$ is
$\mathbf{R}_{\Delta\mathbf{p}}\doteq\mathbb{E}\{(\mathbf{\hat{p}}-\mathbf{p})(\mathbf{\hat{p}}-\mathbf{p}^{H})\}$
$=\mathbf{I}_{n_{B}}-\boldsymbol{\Pi}_{BA}^{\prime}\left(\boldsymbol{\Pi}_{BA}^{\prime 2}+\mathbf{I}_{n_{B}}\right)^{-1}\boldsymbol{\Pi}_{BA}^{\prime}$
$=\left(\boldsymbol{\Pi}_{BA}^{\prime 2}+\mathbf{I}_{n_{B}}\right)^{-1}=\mathbf{I}_{n_{B}}-\mathbf{R}_{\mathbf{\hat{p}}}$ (8)
which is diagonal with the $i$th diagonal element being $\frac{1}{\pi_{BAi}^{\prime 2}+1}=\mathcal{O}(1/p_{A})$.
III-B Phase 2 of G-STEEP
In phase 2, Bob applies Gaussian linear encryption, i.e., he transmits $\sqrt{\frac{p_{B}}{2n_{B}}}\mathbf{x}_{B}=\sqrt{\frac{p_{B}}{2n_{B}}}(\mathbf{\hat{p}}+\mathbf{s})$ where the secret message vector $\mathbf{s}$ is assumed to be $\mathcal{CN}(0,\mathbf{I}_{n_{B}})$. Here $p_{B}$ is the upper bound of the total transmit power from Bob. The corresponding signal received by Alice is
$\mathbf{y}_{A}=\mathbf{H}_{AB}^{\prime\prime}(\mathbf{\hat{p}}+\mathbf{s})+\mathbf{w}_{A}$ (9)
with $\mathbf{H}_{AB}^{\prime\prime}=\sqrt{1/2}\mathbf{H}_{AB}^{\prime}=\sqrt{p_{B}/(2n_{B})}\mathbf{H}_{AB}$. Assume that Alice and Eve both know the feedback of $\mathbf{V}_{BA}$ from Bob. Then, Alice also knows the effective probing vector $\mathbf{p}=\mathbf{V}_{BA}^{H}\mathbf{x}_{A}$.
Note that if $n_{A}=n_{B}=1$, then $\mathbf{V}_{BA}$ would reduce to one and hence the above mentioned feedback would not be needed.
The MMSE estimate of $\mathbf{s}$ by Alice from $\mathbf{y}_{A}$ (and from her knowledge of the exact $\mathbf{x}_{A}$) can be based on this zero-mean sufficient statistic $\Delta\mathbf{y}_{A}\doteq\mathbf{y}_{A}-\mathbb{E}\{\mathbf{y}_{A}|\mathbf{x}_{A}\}$, which can be shown to be
$\Delta\mathbf{y}_{A}=\mathbf{y}_{A}-\mathbf{H}_{AB}^{\prime\prime}\mathbf{R}_{\mathbf{\hat{p}}}\mathbf{p}.$ (10)
Then the MMSE estimate of $\mathbf{s}$ by Alice is
$\mathbf{\hat{s}}_{A}=\mathbf{H}_{AB}^{\prime\prime H}\left(\mathbf{H}_{AB}^{\prime\prime}(\mathbf{R}_{\Delta\mathbf{p}^{\prime}}+\mathbf{I}_{n_{B}})\mathbf{H}_{AB}^{\prime\prime H}+\mathbf{I}_{n_{A}}\right)^{-1}\Delta\mathbf{y}_{A}.$ (11)
Here
$\Delta\mathbf{p}^{\prime}\doteq\mathbf{\hat{p}}-\mathbf{R}_{\mathbf{\hat{p}}}\mathbf{p}=\mathbf{R}_{\Delta\mathbf{p}}\mathbf{\hat{p}}+\mathbf{R}_{\mathbf{\hat{p}}}\Delta\mathbf{p},$ (12)
and
$\mathbf{R}_{\Delta\mathbf{p}^{\prime}}\doteq\mathbb{E}\{\Delta\mathbf{p}^{\prime}\Delta\mathbf{p}^{{}^{\prime}H}\}$
$=\mathbf{R}_{\Delta\mathbf{p}}\mathbf{R}_{\mathbf{\hat{p}}}\mathbf{R}_{\Delta\mathbf{p}}+\mathbf{R}_{\mathbf{\hat{p}}}\mathbf{R}_{\Delta\mathbf{p}}\mathbf{R}_{\mathbf{\hat{p}}}=\mathbf{R}_{\mathbf{\hat{p}}}\mathbf{R}_{\Delta\mathbf{p}}.$ (13)
Then the MSE matrix of $\mathbf{\hat{s}}_{A}$ is
$\mathbf{R}_{\Delta\mathbf{s}_{A}}\doteq\mathbb{E}\{(\mathbf{\hat{s}}_{A}-\mathbf{s}_{A})(\mathbf{\hat{s}}_{A}-\mathbf{s}_{A})^{H}\}$
$=\mathbf{I}_{n_{B}}-\mathbf{H}_{AB}^{\prime\prime H}\left(\mathbf{H}_{AB}^{\prime\prime}(\mathbf{R}_{\Delta\mathbf{p}^{\prime}}+\mathbf{I}_{n_{B}})\mathbf{H}_{AB}^{\prime\prime H}+\mathbf{I}_{n_{A}}\right)^{-1}\mathbf{H}_{AB}^{\prime\prime}$
$=\mathbf{I}_{n_{B}}-\left(\mathbf{H}_{AB}^{\prime\prime H}\mathbf{H}_{AB}^{\prime\prime}(\mathbf{R}_{\Delta\mathbf{p}^{\prime}}+\mathbf{I}_{n_{B}})+\mathbf{I}_{n_{B}}\right)^{-1}\mathbf{H}_{AB}^{\prime\prime H}\mathbf{H}_{AB}^{\prime\prime}$
$=\left(\mathbf{H}_{AB}^{\prime\prime H}\mathbf{H}_{AB}^{\prime\prime}(\mathbf{R}_{\Delta\mathbf{p}^{\prime}}+\mathbf{I}_{n_{B}})+\mathbf{I}_{n_{B}}\right)^{-1}$
$\,\,\cdot\left(\mathbf{H}_{AB}^{\prime\prime H}\mathbf{H}_{AB}^{\prime\prime}\mathbf{R}_{\Delta\mathbf{p}^{\prime}}+\mathbf{I}_{n_{B}}\right).$ (14)
III-C Effective channel capacity from Bob to Alice
The effective channel capacity from Bob to Alice relative to $\mathbf{s}$ (in bits per round-trip symbol interval) is therefore
$C_{A|B,G}\doteq I(\mathbf{s};\{\mathbf{x}_{A},\mathbf{y}_{A}\})=I(\mathbf{s};\mathbf{\hat{s}}_{A})=\log\frac{1}{|\mathbf{R}_{\Delta\mathbf{s}_{A}}|}$
$=\log\frac{|\mathbf{H}_{AB}^{\prime\prime H}\mathbf{H}_{AB}^{\prime\prime}(\mathbf{R}_{\Delta\mathbf{p}^{\prime}}+\mathbf{I}_{n_{B}})+\mathbf{I}_{n_{B}}|}{|\mathbf{H}_{AB}^{\prime\prime H}\mathbf{H}_{AB}^{\prime\prime}\mathbf{R}_{\Delta\mathbf{p}^{\prime}}+\mathbf{I}_{n_{B}}|}\doteq\log\frac{N_{A|B}}{D_{A|B}}$ (15)
where $N_{A|B}$ and $D_{A|B}$ are defined in the obvious way. Here $|A|$ denotes the determinant of $A$. Notice that $\mathbf{s},\mathbf{x}_{A},\mathbf{y}_{A}$ are jointly Gaussian so that the 2nd and 3rd equalities in (III-C) hold. Note that for $p_{A}\to 0\texttt{ or }\infty$, $\mathbf{R}_{\Delta\mathbf{p}^{\prime}}\to\mathbf{0}$ and hence $C_{A|B,G}\to\log|\mathbf{H}_{AB}^{\prime\prime H}\mathbf{H}_{AB}^{\prime\prime}+\mathbf{I}_{n_{B}}|$.
III-D Effective channel capacity from Bob to Eve
After phases 1 and 2 of G-STEEP, the signals received by Eve are
$\mathbf{y}_{EA}=\mathbf{H}_{EA}^{\prime}\mathbf{x}_{A}+\mathbf{w}_{EA},$ (16)
$\mathbf{y}_{EB}=\mathbf{H}_{EB}^{\prime\prime}(\mathbf{\hat{p}}+\mathbf{s})+\mathbf{w}_{EB}.$ (17)
It follows that
$\mathbf{A}\doteq\mathbb{E}\{\mathbf{y}_{EA}\mathbf{y}_{EA}^{H}\}=\mathbf{H}_{EA}^{\prime}\mathbf{H}_{EA}^{\prime H}+\mathbf{I}_{n_{E}},$ (18)
$\mathbf{B}\doteq\mathbb{E}\{\mathbf{y}_{EB}\mathbf{y}_{EB}^{H}\}=\mathbf{H}_{EB}^{\prime\prime}(\mathbf{R}_{\mathbf{\hat{p}}}+\mathbf{I}_{n_{B}})\mathbf{H}_{EB}^{\prime\prime H}+\mathbf{I}_{n_{E}},$ (19)
$\mathbf{C}\doteq\mathbb{E}\{\mathbf{y}_{EA}\mathbf{y}_{EB}^{H}\}=\mathbf{H}_{EA}^{\prime}\mathbf{R}_{\mathbf{x}_{A},\mathbf{\hat{p}}}\mathbf{H}_{EB}^{\prime\prime H},$ (20)
where $\mathbf{R}_{\mathbf{x}_{A},\mathbf{\hat{p}}}\doteq\mathbb{E}\{\mathbf{x}_{A}\mathbf{\hat{p}}^{H}\}$. Let $\mathbf{Q}\doteq[\mathbf{V}_{BA},\mathbf{V}_{BA}^{\perp}]\in\mathcal{C}^{n_{A}\times n_{A}}$ be a unitary matrix. Then
$\mathbf{R}_{\mathbf{x}_{A},\mathbf{\hat{p}}}=\mathbb{E}\{\mathbf{Q}\mathbf{Q}^{H}\mathbf{x}_{A}\mathbf{\hat{p}}^{H}\}=\mathbf{Q}\mathbb{E}\left\{\left[\begin{array}[]{c}\mathbf{V}_{BA}^{H}\mathbf{x}_{A}\mathbf{\hat{p}}^{H}\\ \mathbf{V}_{BA}^{\perp H}\mathbf{x}_{A}\mathbf{\hat{p}}^{H}\end{array}\right]\right\}$ (23)
$=\mathbf{Q}\left[\begin{array}[]{c}\mathbf{R}_{\mathbf{\hat{p}}}\\ 0\end{array}\right]=\mathbf{V}_{BA}\mathbf{R}_{\mathbf{\hat{p}}},$ (26)
where we have used $\mathbb{E}\{\mathbf{p}\mathbf{\hat{p}}^{H}\}=\mathbb{E}\{\mathbf{\hat{p}}\mathbf{\hat{p}}^{H}\}=\mathbf{R}_{\mathbf{\hat{p}}}$ and $\mathbf{V}_{BA}^{\perp H}\mathbb{E}\{\mathbf{x}_{A}\mathbf{\hat{p}}^{H}\}=0$.
The MMSE estimate of $\mathbf{s}$ by Eve from $\mathbf{y}_{EA}$ and $\mathbf{y}_{EB}$ is
$\mathbf{\hat{s}}_{E}=[\mathbf{0}_{n_{B}\times n_{E}},\mathbf{H}_{EB}^{\prime\prime H}]\left[\begin{array}[]{cc}\mathbf{A}&\mathbf{C}\\ \mathbf{C}^{H}&\mathbf{B}\end{array}\right]^{-1}\left[\begin{array}[]{c}\mathbf{y}_{EA}\\ \mathbf{y}_{EB}\end{array}\right]$ (31)
and the MSE matrix of $\mathbf{\hat{s}}_{E}$ is
$\mathbf{R}_{\Delta\mathbf{s}_{E}}\doteq\mathbb{E}\{(\mathbf{\hat{s}}_{E}-\mathbf{s})(\mathbf{\hat{s}}_{E}-\mathbf{s})^{H}\}$
$=\mathbf{I}_{n_{B}}-[\mathbf{0}_{n_{B}\times n_{E}},\mathbf{H}_{EB}^{\prime\prime H}]\left[\begin{array}[]{cc}\mathbf{A}&\mathbf{C}\\ \mathbf{C}^{H}&\mathbf{B}\end{array}\right]^{-1}\left[\begin{array}[]{c}\mathbf{0}_{n_{E}\times n_{B}}\\ \mathbf{H}_{EB}^{\prime\prime}\end{array}\right]$ (36)
$=\mathbf{I}_{n_{B}}-\mathbf{H}_{EB}^{\prime\prime H}(\mathbf{B}-\mathbf{C}^{H}\mathbf{A}^{-1}\mathbf{C})^{-1}\mathbf{H}_{EB}^{\prime\prime}.$ (37)
It follows from (19) and (20) that
$\mathbf{C}^{H}\mathbf{A}^{-1}\mathbf{C}=\mathbf{H}_{EB}^{\prime\prime}\mathbf{T}\mathbf{H}_{EB}^{\prime\prime H}$ (38)
with
$\mathbf{T}=\mathbf{R}_{\mathbf{x}_{A},\mathbf{\hat{p}}}^{H}\mathbf{H}_{EA}^{\prime H}\left(\mathbf{H}_{EA}^{\prime}\mathbf{H}_{EA}^{\prime H}+\mathbf{I}_{n_{E}}\right)^{-1}\mathbf{H}_{EA}^{\prime}\mathbf{R}_{\mathbf{x}_{A},\mathbf{\hat{p}}}$
$=\mathbf{R}_{\mathbf{\hat{p}}}^{H}\mathbf{V}_{BA}^{H}\left(\mathbf{H}_{EA}^{\prime H}\mathbf{H}_{EA}^{\prime}+\mathbf{I}_{n_{A}}\right)^{-1}\mathbf{H}_{EA}^{\prime H}\mathbf{H}_{EA}^{\prime}\mathbf{V}_{BA}\mathbf{R}_{\mathbf{\hat{p}}}.$ (39)
Let
$\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}\doteq\mathbf{R}_{\mathbf{\hat{p}}}-\mathbf{T}$ (40)
which we see is the MSE matrix of the MMSE estimate $\mathbf{\hat{p}}_{E}$ of $\mathbf{\hat{p}}$ from $\mathbf{y}_{EA}$. Hence, $\mathbf{T}=\mathbf{R}_{\mathbf{\hat{p}}_{E}}\doteq\mathbb{E}\{\mathbf{\hat{p}}_{E}\mathbf{\hat{p}}_{E}^{H}\}$.
Then (III-D) becomes
$\mathbf{R}_{\Delta\mathbf{s}_{E}}=\mathbf{I}_{n_{B}}-\mathbf{H}_{EB}^{\prime\prime H}$
$\,\,\cdot\left(\mathbf{H}_{EB}^{\prime\prime}(\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}+\mathbf{I}_{n_{B}})\mathbf{H}_{EB}^{\prime\prime H}+\mathbf{I}_{n_{E}}\right)^{-1}\mathbf{H}_{EB}^{\prime\prime}$
$=\mathbf{I}_{n_{B}}-\left(\mathbf{H}_{EB}^{\prime\prime H}\mathbf{H}_{EB}^{\prime\prime}(\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}+\mathbf{I}_{n_{B}})+\mathbf{I}_{n_{B}}\right)^{-1}\mathbf{H}_{EB}^{\prime\prime H}\mathbf{H}_{EB}^{\prime\prime}$
$=\left(\mathbf{H}_{EB}^{\prime\prime H}\mathbf{H}_{EB}^{\prime\prime}(\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}+\mathbf{I}_{n_{B}})+\mathbf{I}_{n_{B}}\right)^{-1}$
$\,\,\cdot\left(\mathbf{H}_{EB}^{\prime\prime H}\mathbf{H}_{EB}^{\prime\prime}\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}+\mathbf{I}_{n_{B}}\right).$ (41)
Hence the capacity of the effective return channel from Bob to Eve relative to $\mathbf{s}$ (in bits per round-trip symbol interval) is
$C_{E|B,G}\doteq I(\mathbf{s};\{\mathbf{y}_{EA},\mathbf{y}_{EB}\})=I(\mathbf{s};\mathbf{\hat{s}}_{E})=\log\frac{1}{|\mathbf{R}_{\Delta\mathbf{s}_{E}}|}$
$=\log\frac{|\mathbf{H}_{EB}^{\prime\prime H}\mathbf{H}_{EB}^{\prime\prime}(\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}+\mathbf{I}_{n_{B}})+\mathbf{I}_{n_{B}}|}{|\mathbf{H}_{EB}^{\prime\prime H}\mathbf{H}_{EB}^{\prime\prime}\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}+\mathbf{I}_{n_{B}}|}$
$\doteq\log\frac{N_{E|B}}{D_{E|B}}.$ (42)
Again we have applied the jointly Gaussian nature of $\mathbf{s},\mathbf{y}_{EA},\mathbf{y}_{EB}$ for the 2nd and 3rd equalities in (III-D).
III-E Secrecy capacity of G-STEEP
Theorem 1
An achievable secrecy rate of G-STEEP based on the effective wiretap-channel system from Bob to Alice against Eve (in bits per round-trip symbol interval) is
$C_{s,G}\doteq(I(\mathbf{s};\{\mathbf{x}_{A},\mathbf{y}_{A}\})-I(\mathbf{s};\{\mathbf{y}_{EA},\mathbf{y}_{EB}\}))^{+}$
$=\left(C_{A|B,G}-C_{E|B,G}\right)^{+}=\left(\log\frac{N_{A|B}D_{E|B}}{D_{A|B}N_{E|B}}\right)^{+}$
$=\left[\log\left(\frac{|\mathbf{H}_{AB}^{\prime\prime H}\mathbf{H}_{AB}^{\prime\prime}(\mathbf{R}_{\Delta\mathbf{p}^{\prime}}+\mathbf{I}_{n_{B}})+\mathbf{I}_{n_{B}}|}{|\mathbf{H}_{AB}^{\prime\prime H}\mathbf{H}_{AB}^{\prime\prime}\mathbf{R}_{\Delta\mathbf{p}^{\prime}}+\mathbf{I}_{n_{B}}|}\right.\right.$
$\,\,\cdot\left.\left.\frac{|\mathbf{H}_{EB}^{\prime\prime H}\mathbf{H}_{EB}^{\prime\prime}\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}+\mathbf{I}_{n_{B}}|}{|\mathbf{H}_{EB}^{\prime\prime H}\mathbf{H}_{EB}^{\prime\prime}(\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}+\mathbf{I}_{n_{B}})+\mathbf{I}_{n_{B}}|}\right)\right]^{+}.$ (43)
where $\mathbf{R}_{\Delta\mathbf{p}^{\prime}}$ is given in (III-B), and $\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}$ is given in (40).
Proof:
This follows from the WTC theory for Gaussian channels [2] with respect to the message $\mathbf{s}$ from Bob, and the previous results shown in (III-C) and (III-D). ∎
One may argue that the secret-key capacity based on $\{\mathbf{x}_{A},\mathbf{y}_{A}\}$ at Alice, $\{\mathbf{y}_{B},\mathbf{s}\}$ at Bob and $\{\mathbf{y}_{EA},\mathbf{y}_{EB}\}$ at Eve after both phases of G-STEEP (and using additional and iterative operations for information reconciliation and privacy amplification via public communications) should be larger than or equal to $C_{s,G}$. But this paper will not pursue that direction of exploration because WTC transmission is relatively simple and its secrecy rate over the effective WTC system induced by STEEP is already promising as discussed in the rest of this section.
Corollary 1
If $n_{B}=1$ and $\mathbf{H}_{AB}$, $\mathbf{H}_{BA}$ and $\mathbf{H}_{EB}$ are replaced by $\mathbf{h}_{AB}$, $\mathbf{h}_{BA}$ and $\mathbf{h}_{EB}$ (similarly for their scaled versions), then $C_{s,G}\doteq\left(C_{A|B,G}-C_{E|B,G}\right)^{+}$ with
$C_{A|B,G}=\log\left(1+\frac{\frac{S_{AB}}{2}}{\frac{1}{2}\frac{S_{AB}S_{BA}}{(S_{BA}+1)^{2}}+1}\right),$ (44)
$C_{E|B,G}=\log\left(1+\frac{\frac{S_{EB}}{2}}{(\sigma_{\hat{p}_{0}}^{2}-t)\frac{S_{EB}}{2}+1}\right),$ (45)
where $S_{AB}=\|\mathbf{h}_{AB}^{\prime}\|^{2}$, $S_{BA}=\|\mathbf{h}_{BA}^{\prime}\|^{2}$, $S_{EB}=\|\mathbf{h}_{EB}^{\prime}\|^{2}$, $\sigma^{2}_{\hat{p}_{0}}=\frac{S_{BA}}{S_{BA}+1}$ and
$t=\mathbf{r}^{H}\mathbf{H}_{EA}^{\prime H}\left(\mathbf{H}_{EA}^{\prime}\mathbf{H}_{EA}^{\prime H}+\mathbf{I}_{n_{E}}\right)^{-1}\mathbf{H}_{EA}^{\prime}\mathbf{r}$ (46)
with $\mathbf{r}=\sigma^{2}_{\hat{p}_{0}}\frac{1}{\|\mathbf{h}_{BA}\|}\mathbf{h}_{BA}^{*}$.
Proof:
This follows from Theorem 1. In particular, $\mathbf{T}$ in (III-D) is now reduced to the scalar $t$. ∎
We see that for the case of $n_{B}=1$, the effects of $\mathbf{h}_{AB}^{\prime}$, $\mathbf{h}_{BA}^{\prime}$ and $\mathbf{h}_{EB}^{\prime}$ on $C_{s,G}$ are only through their norms. The effect of $\mathbf{H}_{EA}^{\prime}$ on $C_{s,G}$ is only through the scalar $t$.
III-F Properties of $C_{s,G}$
Lemma 1
Assuming constant channel matrices, the secret-key capacity $C_{key}$ (in bits per probing symbol interval) based on the data sets at Alice, Bob and Eve after phase 1 of G-STEEP (and a public communication phase after that) is
$C_{key}=\log\left|\mathbf{I}_{n_{A}}+\mathbf{H}_{BA}^{\prime H}\mathbf{H}_{BA}^{\prime}\left(\mathbf{H}_{EA}^{\prime H}\mathbf{H}_{EA}^{\prime}+\mathbf{I}_{n_{A}}\right)^{-1}\right|.$ (47)
Proof:
This follows from Theorem 1 in [11] where Maurer’s lower and upper bounds, generalized in [2] and further applied asymptotically to continuous sources via generalized mutual information, are used. (Some earlier attempt such as [26] to improve Maurer’s lower bound is not necessary here.) $C_{key}$ here is $\xi_{B}$ in equation (8) in [11] with constant channel matrices, and $\mathbf{H}_{EA}$ and $\frac{p_{A}}{n_{A}}$ here are $\mathbf{G}_{A}$ and $\gamma_{BA}$ in [11]. Furthermore, $\lambda_{B}$ and $\lambda_{EA}$ in [11] are both normalized to be one here. ∎
Note that $C_{key}$ is the maximum secret-key rate achievable based on the data sets generated in phase 1 of G-STEEP at Alice, Bob and Eve and through communications in the public network (where Eve has access to all communications). So, if Eve’s receive channel from Bob is no weaker than Alice’s receive channel from Bob (in phase 2 of G-STEEP), we should expect $C_{s,G}\leq C_{key}$. If $C_{s,G}$ approaches $C_{key}$ under high powers, we can say that $C_{s,G}$ is optimal against strong Eve under high powers.
But if Eve’s receive channel from Bob is weaker than that at Alice, it is possible to have $C_{s,G}>C_{key}$. But one should not be excited by this situation. We know that $C_{key}$ is based on the assumption that all communications for secret key generation are done in the public domain. If any of these communications are not public, the resulting $C_{key}$ would be higher. So, for a meaningful comparison between $C_{s,G}$ and $C_{key}$, we should assume that Eve’s receive channel in phase 2 of G-STEEP is no weaker than that at Alice.
Proposition 1
Assume that $\mathbf{H}_{AB}$, $\mathbf{H}_{BA}$, $\mathbf{H}_{EA}$ and $\mathbf{H}_{EB}$ are typical realizations (where the rank of each matrix equals to the minimum of its numbers of rows and columns and the rank conditions in (-A), (123) and (129) hold). For $n_{A}\geq n_{B}$, $n_{E}\geq 1$ and any given (fixed) $\eta_{p}=\frac{p_{B}}{p_{A}}$,
$\lim_{p_{A}\to\infty}\frac{1}{\log p_{A}}C_{s,G}=\lim_{p_{A}\to\infty}\frac{1}{\log p_{A}}C_{key}$
$=\min(n_{B},(n_{A}-n_{E})^{+}),$ (48)
i.e., $\texttt{DoF}(C_{s,G})=\texttt{DoF}(C_{key})$. Namely, $C_{s,G}$ is optimal in DoF.
Proof:
See Appendix-A. ∎
The above DoF is the maximum achievable DoF currently known, which is consistent with a prior result in [12].
The DoF only depends on the numbers of antennas on Alice, Bob and Eve, which is not affected by any finite scaling on channel matrices and/or on noise variances.
Proposition 2
Assume typical realizations of all channel matrices (like those in Proposition 1). For $n_{E}\geq n_{A}\geq n_{B}\geq 1$,
$\lim_{p_{A}\to\infty}(\lim_{p_{B}\to\infty}C_{s,G})=\lim_{p_{A}\to\infty}C_{key}$
$=\log\left|\mathbf{I}_{n_{B}}+\boldsymbol{\Pi}_{BA}^{2}\mathbf{V}_{BA}^{H}(\mathbf{H}_{EA}^{H}\mathbf{H}_{EA})^{-1}\mathbf{V}_{BA}\right|.$ (49)
Namely, $C_{s,G}$ is optimal (against strong Eve) asymptotically as $p_{A}\to\infty$ and $\frac{p_{B}}{p_{A}}\to\infty$.
Proof:
See Appendix-B. ∎
The above proposition is also intuitively justified if we think of $\frac{p_{B}}{p_{A}}\to\infty$ as somewhat similar to the case where phase 2 of G-STEEP only uses public communications and also think of $p_{A}\to\infty$ as somewhat similar to the case where the encryption in phase 2 is done via a modulo sum between two discrete random variables. In other words, for $\frac{p_{B}}{p_{A}}\to\infty$, both Alice and Eve would receive the same $\sqrt{p_{B}}(\mathbf{\hat{p}}+\mathbf{s})$ from Bob, i.e., the phase 2 would be via a public channel. For $p_{B}\to\infty$, $\sqrt{p_{B}}(\mathbf{\hat{p}}+\mathbf{s})=\sqrt{p_{B}}\mathbf{\hat{p}}+\sqrt{p_{B}}\mathbf{s}$ is a sum between $\sqrt{p_{B}}\mathbf{\hat{p}}$ and $\sqrt{p_{B}}\mathbf{s}$ (which would be virtually uniformly distributed), and this sum would be like a modulo-sum with an infinite modulo. Then the encryption lemma would suggest that in the case of $p_{A}\to\infty$ and $\frac{p_{B}}{p_{A}}\to\infty$, $C_{key}=C_{s,G}$. Again, the above discussion is no proof but only an intuition to make an intuitive sense of the result actually proven in Appendix-B.
Since the limit in (2) is always positive (unless $\mathbf{H}_{EA}$ has an infinite norm), this proposition also suggests that for a sufficiently large (but finite) $p_{A}$ and a sufficiently large (but finite) $\frac{p_{B}}{p_{A}}$, $C_{s,G}$ is positive. We will see a more specific case of this next.
III-G The special case of $n_{A}=n_{B}=1$
For $n_{A}=n_{B}=1$, we let $\mathbf{H}_{AB}$, $\mathbf{H}_{BA}$, $\mathbf{H}_{EA}$ and $\mathbf{H}_{EB}$ be replaced by $h_{AB}$, $h_{BA}$, $\mathbf{h}_{EA}$ and $\mathbf{h}_{EB}$. Then it follows from Corollary 1 that
$C_{s,G}=\left[\log\left(1+\frac{b/2}{bA_{1}/2+1}\right)-\log\left(1+\frac{\beta b/2}{\beta bA_{2}/2+1}\right)\right]^{+}$ (50)
with $a\doteq S_{BA}\doteq p_{A}|h_{BA}|^{2}=|h_{BA}^{\prime}|^{2}$ and $b\doteq S_{AB}\doteq p_{B}|h_{AB}|^{2}=|h_{AB}^{\prime}|^{2}$. Also $A_{1}=\frac{a}{(a+1)^{2}}$ and $A_{2}=A_{1}\frac{(a+\alpha a+1)}{\alpha a+1}$ with $\alpha\doteq\frac{S_{EA}}{S_{BA}}$ and $\beta\doteq\frac{S_{EB}}{S_{AB}}$. Furthermore, $S_{EA}\doteq p_{A}\|\mathbf{h}_{EA}\|^{2}=\|\mathbf{h}_{EA}^{\prime}\|^{2}$ and $S_{EB}\doteq p_{B}\|\mathbf{h}_{EB}\|^{2}=\|\mathbf{h}_{EB}^{\prime}\|^{2}$. Note that $A_{2}>A_{1}$ and they are invariant to $b$.
In this special case, all channel gains and noise variances are completely lumped into just four parameters: $a$, $b$, $\alpha$ and $\beta$. Here $a$ and $b$ are respectively the (raw channel) SNR at Bob in phase 1 and the (raw channel) SNR at Alice in phase 2. And $a$ and $b$ are proportional to $p_{A}$ and $p_{B}$ respectively. Furthermore, $\alpha$ and $\beta$ are the SNR ratios measuring Eve’s (raw) channel strengths over users’ (raw) channel strengths in phases 1 and 2 respectively. It is important to distinguish between “raw channels” and “effective channels”, the latter of which are induced by STEEP.
In particular, if Eve’s (raw) channel is stronger than users’ (raw) channel in phase 1, then $\alpha>1$; and if Eve’s (raw) channel is stronger than users’ (raw) channel in phase 2, then $\beta>1$. It is obvious that $C_{s,G}$ increases as $\alpha$ and/or $\beta$ decrease.
If $\alpha\geq 1$ and $\beta\geq 1$, all conventional WTC schemes either from Alice to Bob or from Bob to Alice have zero secrecy capacity. (In the MIMO case, a similar conclusion can be drawn if $n_{E}\geq n_{A}\geq n_{B}$ and the large-scale fading at Eve is smaller than those at Alice and Bob.)
Proposition 3
Assume $n_{A}=n_{B}=1$. Then $C_{s,G}>0$ if and only if
$b>\bar{b}\doteq\frac{2(\beta-1)}{\beta(A_{2}-A_{1})}$
$=\frac{2(\beta-1)(a+1)^{2}(\alpha a+1)}{\beta a^{2}}.$ (51)
Also for $\beta>1$, $C_{s,G}$ increases as $b$ increases.
Proof:
This can be directly verified from (50). ∎
We see that as $a$ (or equivalently $p_{A}$) either decreases to zero or increases to infinity, $\bar{b}$ increases to infinity subject to $\beta>1$. But for $\alpha>1$, $\beta>1$ and $a\gg 1$, we have
$\bar{b}\approx 2\frac{\beta-1}{\beta}\alpha a=\mathcal{O}(a)$ (52)
Hence, for $\alpha>1$, $\beta>1$ and a given $b$ (such as 20dB to 30dB in practice), the optimal value of $a$ to maximize $C_{s,G}$ is generally in between zero and $b$.
In practice, we can utilize (3) to ensure a positive secrecy rate whenever an upper bound on $\alpha$ is available. In the case of random fading channels, the probability for (3) not to hold can be kept small by keeping a large ratio of $p_{B}$ over $p_{A}$ [29].
III-G1 Comparison to $C_{key}$
For $n_{A}=n_{B}=1$, (47) reduces to
$C_{key}=\log\left(1+\frac{S_{BA}}{S_{EA}+1}\right)=\log\left(1+\frac{a}{\alpha a+1}\right)$
$=\log\frac{A_{2}}{A_{1}}.$ (53)
It follows from (50) that
$\lim_{b\to\infty}C_{s,G}=\left(\log\left(1+\frac{1}{A_{1}}\right)-\log\left(1+\frac{1}{A_{2}}\right)\right)^{+}$
$=\left(\log\frac{A_{2}(A_{1}+1)}{A_{1}(A_{2}+1)}\right)^{+}<C_{key}$ (54)
Since $C_{s,G}$ increases with $b$ for $\beta>1$, then for $\beta>1$ we have $C_{s,G}<C_{key}$ for all $\alpha$, $a$ and $b$. This is expected as discussed before.
However,
$\lim_{a\to\infty}C_{key}=\log\left(1+\frac{1}{\alpha}\right),$ (55)
which is the same as
$\lim_{a\to\infty}(\lim_{b\to\infty}C_{s,G})$
$=\lim_{a\to\infty}\left(\log\frac{A_{2}(A_{1}+1)}{A_{1}(A_{2}+1)}\right)^{+}$
$=\log\left(1+\frac{1}{\alpha}\right).$ (56)
So, if both $a$ and $b$ are large but $b$ dominates $a$, then $C_{s,G}=C_{key}$. This is a special case of Proposition 2.
IV STEEP with PSK channel Probing and PSK Nonlinear Encryption (P-STEEP)
In this section, P-STEEP is presented assuming $n_{A}=n_{B}=1$ and $n_{E}\geq 1$. It is important to note that for applications where power control is difficult (due to nonlinearity of power amplifier, channel disturbances, etc), nonlinear modulation such as PSK is always preferred to linear modulation.
IV-A Operations at Alice and Bob
In phase 1 of P-STEEP, Alice sends out PSK probes $\sqrt{p_{A}}x_{A}=\sqrt{p_{A}}e^{j\theta}$ where $\theta$ is an M-ary discrete uniform random variable within $[-\pi,\pi]$. Then Bob receives
$y_{B}=\sqrt{p_{A}}h_{BA}x_{A}+w_{B}.$ (57)
A sufficient statistic from $y_{B}$ for $x_{A}=e^{j\theta}$ (at Bob) is
$r_{B}\doteq\frac{1}{\sqrt{p_{A}}h_{BA}}y_{B}=x_{A}+v_{B}$ (58)
where $v_{B}$ is $\mathcal{CN}(0,\frac{1}{S_{BA}})$ with $S_{BA}=p_{A}|h_{BA}|^{2}$.
In phase 2 of P-STEEP, Bob applies PSK nonlinear encryption, i.e., he sends out $\sqrt{p_{B}}x_{B}=\sqrt{p_{B}}e^{j\phi}r_{B}$ where $\phi$ is a secret phase value (meant for Alice) randomly chosen from the same discrete constellation as $\theta$. Here the construction of $x_{B}=e^{j\phi}r_{B}$ is different from that for G-STEEP with $n_{A}=n_{B}=1$. This nonlinear encryption fits naturally with PSK (a nonlinear modulation).
It is important to note that while both $\theta$ and $\phi$ are discrete, $r_{B}$ here is continuous. The use of continuous $r_{B}$ (instead of a quantized $r_{B}$ with the constellation size $M$) to construct $x_{B}=e^{j\phi}r_{B}$ reduces the computational complexity at Bob (i.e., no detection is needed at Bob). It is however not clear whether this would yield a better secrecy rate than the quantized option. There is also a strategy in between “completely hard” and “completely soft”, i.e., replacing $r_{B}$ by its quantized value with a constellation size equal to $lM$ with $l\geq 1$. When $l=1$, we say that the quantized $r_{B}$ is completely hard. As $l$ becomes larger, the quantized $r_{B}$ becomes “softer”. But in this paper, we only focus on continuous $r_{B}$.
Then Alice receives
$y_{A}=\sqrt{p_{B}}h_{AB}x_{B}+w_{A}.$ (59)
A sufficient statistic from $y_{A}$ for $\phi$ (at Alice) is
$r_{A}\doteq\frac{x_{A}^{*}}{\sqrt{p_{B}}h_{AB}}y_{A}=e^{j\phi}+e^{j\phi}x_{A}^{*}v_{B}+x_{A}^{*}v_{A}$ (60)
where $v_{A}$ is $\mathcal{CN}(0,\frac{1}{S_{AB}})$ with $S_{AB}=p_{B}|h_{AB}|^{2}$.
Since $v_{B}$ and $v_{A}$ are independent circular complex Gaussian, we can also write
$r_{A}\doteq\frac{x_{A}^{*}}{\sqrt{p_{B}}h_{AB}}y_{A}=e^{j\phi}+v_{B}^{\prime}+v_{A}^{\prime}$ (61)
where $v_{B}^{\prime}$ and $v_{A}^{\prime}$ are independent of $\phi$ and each other, and they have the same distributions as $v_{B}$ and $v_{A}$.
The minimum distance between the constellation points of $e^{j\phi}$ is $2\sin\frac{\pi}{M}$. Hence the error rate in detecting $e^{j\phi}$ from $r_{A}$ is (approximately for $M=2^{m}$ with $m\geq 2$)
$p_{e,A}=n_{0}Q\left(\frac{\sin\frac{\pi}{M}}{\epsilon_{A}}\right)$ (62)
where $n_{0}=1$ for $m=1$, $n_{0}=2$ for $m\geq 2$,
$\epsilon_{A}=\sqrt{\frac{1}{2S_{BA}}+\frac{1}{2S_{AB}}}$ (63)
[…]
The effective capacity from Bob to Alice relative to $\phi$ is
$C_{A|B,P}\doteq I(\phi;r_{A})=H(\phi)-H(\phi|r_{A}),$ (64)
where $H(\phi)=\log M$ (the entropy of $\phi$). To determine $H(\phi|r_{A})$, we can view $\phi$ given $r_{A}$ as the optimal decision of $\phi$ from $r_{A}$.
For $M=2$, the optimal decision of $\phi$ from $r_{A}$ takes two possible values with the probabilities $1-p_{e,A}$ and $p_{e,A}$ respectively. In this case,
$C_{A|B,P}=1-h_{2}(p_{e,A})$ (65)
with $h_{2}(p)\doteq-p\log p-(1-p)\log(1-p)$.
For $M=2^{m}$ with $m\geq 2$, the optimal decision of $\phi$ from $r_{A}$ takes approximately three possible values with the probabilities $1-p_{e,A}$, $\frac{1}{2}p_{e,A}$ and $\frac{1}{2}p_{e,A}$ respectively. In this case, we can write
$C_{A|B,P}=m-p_{e,A}-h_{2}(p_{e,A})$ (66)
with $m\geq 2$.
IV-B Operations at Eve
After phases 1 and 2, the signals received by Eve are
$\mathbf{y}_{EA}=\sqrt{p_{A}}\mathbf{h}_{EA}x_{A}+\mathbf{w}_{EA},$ (67)
$\mathbf{y}_{EB}=\sqrt{p_{B}}\mathbf{h}_{EB}x_{B}+\mathbf{w}_{EB},$ (68)
or equivalently
$r_{EA}\doteq\frac{\mathbf{h}_{EA}^{H}}{\sqrt{p_{A}}\|\mathbf{h}_{EA}\|^{2}}\mathbf{y}_{EA}=x_{A}+v_{EA},$ (69)
$r_{EB}\doteq\frac{\mathbf{h}_{EB}^{H}}{\sqrt{p_{A}}\|\mathbf{h}_{EB}\|^{2}}\mathbf{y}_{EB}=x_{B}+v_{EB}.$ (70)
Here $v_{EA}$ is $\mathcal{CN}(0,\frac{1}{S_{EA}})$ with $S_{EA}=p_{A}\|\mathbf{h}_{EA}\|^{2}$, and $v_{EB}$ is $\mathcal{CN}(0,\frac{1}{S_{EB}})$ with $S_{EB}=p_{B}\|\mathbf{h}_{EB}\|^{2}$.
Let us consider
$r_{E}\doteq r_{EA}^{*}r_{EB}=e^{j\phi}+x_{A}^{*}e^{j\phi}v_{B}+x_{A}^{*}v_{EB}+v_{EA}^{*}e^{j\phi}x_{A}$ (71)
where we have ignored the second-order terms of noises: $e^{j\phi}v_{EA}^{*}v_{B}$ and $v_{EA}^{*}v_{EB}$. Since $v_{B}$, $v_{EB}$ and $v_{EA}$ are independent circular complex Gaussian, we can also write
$r_{E}\doteq r_{EA}^{*}r_{EB}=e^{j\phi}+v_{B}^{\prime}+v_{EB}^{\prime}+v_{EA}^{\prime}$ (72)
where $v_{B}^{\prime}$, $v_{EB}^{\prime}$ and $v_{EA}^{\prime}$ are also independent circular complex Gaussian and are independent of $\phi$ and $x_{A}$, and they have the same distributions as $v_{B}$, $v_{EB}$ and $v_{EA}$ respectively.
Since $\{r_{EA},r_{EB}\}$ is a one-to-one function of $\{r_{EA},r_{E}\}$, and $r_{EA}$ is independent of $r_{E}$ and $\phi$, we now know that $r_{E}$ is a sufficient statistic from $\{r_{EA},r_{EB}\}$ for $\phi$.
So the optimal detection of $e^{j\phi}$ from $\{r_{EA},r_{EB}\}$ is the same as that from $r_{E}$. We know that the error rate in detecting $e^{j\phi}$ from $r_{E}$ is (approximately for $M=2^{m}\geq 4$)
$p_{e,E}=n_{0}Q\left(\frac{\sin\frac{\pi}{M}}{\epsilon_{E}}\right)$ (73)
where
$\epsilon_{E}=\sqrt{\frac{1}{2S_{BA}}+\frac{1}{2S_{EA}}+\frac{1}{2S_{EB}}}.$ (74)
Similar to $C_{A|B,P}$ in (65) and (66), we can express the effective capacity $C_{E|B,P}$ from Bob to Eve relative to $\phi$ as
$C_{E|B,P}=m-p_{e,E}-h_{2}(p_{e,E}).$ (75)
IV-C Secrecy capacity
The secrecy capacity of P-STEEP is
$C_{s,P}\doteq(C_{A|B,P}-C_{E|B,P})^{+}=h_{2}(p_{e,E})-h_{2}(p_{e,A})$ (76)
which is positive if and only if $p_{e,A}<p_{e,E}$. We see that $p_{e,A}<p_{e,E}$ if and only if $\epsilon_{A}<\epsilon_{E}$. It follows from (63) and (74) that
$\frac{\epsilon_{A}^{2}}{\epsilon_{E}^{2}}=\frac{\frac{1}{a}+\frac{1}{b}}{\left(1+\frac{1}{\alpha}\right)\frac{1}{a}+\frac{1}{\beta b}},$ (77)
where $a=S_{BA}$, $b=S_{AB}$, $\alpha=\frac{S_{EA}}{S_{BA}}$ and $\beta=\frac{S_{EB}}{S_{AB}}$. Hence $\epsilon_{A}<\epsilon_{E}$ if and only if
$\frac{b}{a}>\alpha\left(1-\frac{1}{\beta}\right).$ (78)
The condition (78) always holds if $\beta<1$ (i.e., Eve’s receive channel from Bob is weaker than Alice’s receive channel from Bob). Otherwise, for $\beta>1$, the condition (78) can be met by a sufficiently large but finite $p_{B}$ while $p_{A}$ is finite (subject to all other parameters being finite).
IV-D Ratio of bit error rates
For large $a$ and $b$, both $p_{e,A}$ and $p_{e,E}$ are small subject to $\alpha>1$ and $\beta>1$. In this case, $C_{s,P}$ is only a small positive value under (78). But the ratio $\gamma_{p}$ of $p_{e,A}$ over $p_{e,E}$ is also a meaningful metrics subject to a sufficiently long packet (e.g., a packet of $n$ independent bits with $np_{e,E}\approx 1$).
Applying $\frac{x}{1+x^{2}}\phi_{0}(x)<Q(x)<\frac{\phi_{0}(x)}{x}$ with $\phi_{0}(x)=\frac{1}{\sqrt{2\pi}}e^{-\frac{x^{2}}{2}}$ and the condition (78), one can verify that
$\gamma_{p}\doteq\frac{p_{e,A}}{p_{e,E}}<(1+\delta_{p})\exp\left(-P\right)$ (79)
where
$\delta_{p}=\frac{\epsilon_{A}\epsilon_{E}}{\sin^{2}\frac{\pi}{M}},$ (80)
$P=\sin^{2}\left(\frac{\pi}{M}\right)\frac{a^{2}b(\beta b+\alpha a-\alpha\beta a)}{(a+b)(\alpha\beta ab+\beta ab+\alpha a^{2})}.$ (81)
Here $1+\delta_{p}\approx 1$ for large $a$ and $b$. To obtain a large $P$ and hence a very small $\gamma_{p}$, we need a large $a$ and a large $\frac{b}{a}$ because
$\lim_{b\to\infty}P=\sin^{2}\left(\frac{\pi}{M}\right)\frac{a}{\alpha+1}$ (82)
which increases with $a$. For example, if $M=2$, $\alpha=\beta=2$, $a=10^{2}$ and $b=10^{3}$ (i.e., 20dB and 30dB respectively), we have $P\approx 26.4$.
V STEEP for Multiple Access (M-STEEP)
Let us now go back to G-STEEP but consider its use for multiple access. Specifically, let there be an access point (AP) with $n_{A}$ antennas, and $M$ units of single-antenna user equipment (UE) which are denoted by UE1, $\cdots$, UEM. If we apply G-STEEP to AP and each UE separately, there would be a significant overhead associated with the channel probing for each UE. To reduce the overhead, an option is to allow all UEs to take advantage of the same probes transmitted by the AP. We will show a power condition under which the secrecy rate from each UE to AP stays positive for any given $M$.
In phase 1 of M-STEEP, AP broadcasts a sequence of independent realizations of the random probing vector $\sqrt{p_{A}/n_{A}}\mathbf{x}\in\mathcal{C}^{n_{A}\times 1}$ with $\mathbf{x}$ being $\mathcal{CN}(\mathbf{0},\mathbf{I})$. Then UEi receives
$y_{i}=\sqrt{p_{A}/n_{A}}\mathbf{h}_{i}^{T}\mathbf{x}+w_{i}=\mathbf{h}_{i}^{\prime T}\mathbf{x}+w_{i}$ (83)
with $i=1,\cdots,M$, $\mathbf{h}_{i}^{\prime}=\sqrt{p_{A}/n_{A}}\mathbf{h}_{i}$ and $w_{i}$ being $\mathcal{CN}(0,1)$. The effective probe arriving at UEi is defined to be $p_{i}=\mathbf{\bar{h}}_{i}^{T}\mathbf{x}$ with $\mathbf{\bar{h}}_{i}^{T}\doteq\frac{\mathbf{h}_{i}^{T}}{\|\mathbf{h}_{i}\|}$. The MMSE estimate of $p_{i}$ is denoted by $\hat{p}_{i}$, and its MSE is
$d_{i}\doteq\frac{1}{S_{i}+1}$ (84)
with $S_{i}=(p_{A}/n_{A})\|\mathbf{h}_{i}\|^{2}$. The variance of $\hat{p}_{i}$ is
$c_{i}\doteq 1-d_{i}=\frac{S_{i}}{S_{i}+1}.$ (85)
One can also verify that $\mathbb{E}\{p_{i}p_{j}^{*}\}=\phi_{i,j}$, $\mathbb{E}\{\hat{p}_{i}\hat{p}_{j}^{*}\}=c_{i}c_{j}\phi_{i,j}$, and $\mathbb{E}\{p_{i}\hat{p}_{j}^{*}\}=c_{j}\phi_{i,j}$ with $\phi_{i,j}=\mathbf{\bar{h}}_{i}^{T}\mathbf{\bar{h}}_{j}^{*}$.
In phase 2 of M-STEEP, the UEs use orthogonal multiple access to the AP. Specifically, UEi sends out a sequence of random realizations of $\sqrt{p_{ui}/2}(\hat{p}_{i}+s_{i})$ (of power upper bounded by $p_{ui}$) with $s_{i}$ being a secret random symbol with the distribution $\mathcal{CN}(0,1)$, and the corresponding signal received by the AP is
$\mathbf{y}_{Ai}=\sqrt{p_{ui}/2}(\hat{p}_{i}+s_{i})\mathbf{h}_{Ai}+\mathbf{w}_{Ai}$
$=\sqrt{1/2}(\hat{p}_{i}+s_{i})\mathbf{h}_{Ai}^{\prime}+\mathbf{w}_{Ai}\in\mathcal{C}^{n_{A}\times 1}$ (86)
with $\mathbf{w}_{Ai}$ being $\mathcal{CN}(\mathbf{0},\mathbf{I})$ and $\mathbf{h}_{Ai}^{\prime}=\sqrt{p_{ui}}\mathbf{h}_{Ai}$.
It follows from (III-C) with $n_{B}=1$ or from (44) that the effective capacity from UEi to AP relative to $s_{i}$ is
$C_{A|i}\doteq I(s_{i};\hat{s}_{i})=\log\left(1+\frac{\frac{S_{Ai}}{2}}{\frac{S_{i}S_{Ai}/2}{(S_{i}+1)^{2}}+1}\right).$ (87)
where $\hat{s}_{i}$ is the MMSE estimate of $s_{i}$ by AP, $S_{Ai}=p_{ui}\|\mathbf{h}_{Ai}\|^{2}$ and $S_{i}$ was defined before.
V-A Effective Return Channel from each UE to Eve
The signals received by Eve during both phases of M-STEEP are
$\mathbf{y}_{EA}=\sqrt{p_{A}/n_{A}}\mathbf{H}_{EA}\mathbf{x}+\mathbf{w}_{EA}=\mathbf{H}_{EA}^{\prime}\mathbf{x}+\mathbf{w}_{EA},$ (88)
$\mathbf{y}_{Ei}=\sqrt{p_{ui}/2}\mathbf{h}_{Ei}(\hat{p}_{i}+s_{i})+\mathbf{w}_{Ei}$
$=\sqrt{1/2}\mathbf{h}_{Ei}^{\prime}(\hat{p}_{i}+s_{i})+\mathbf{w}_{Ei},$ (89)
for all $i=1,\cdots,M$. Here $\mathbf{h}_{Ei}^{\prime}=\sqrt{p_{ui}}\mathbf{h}_{Ei}$, and $\hat{p}_{i}$ for every $i$ depends on $\mathbf{x}$. Also note that $s_{1},\cdots,s_{M}$ are independent of each other.
It can be shown that the MSE of the MMSE estimate $\hat{s}_{iE}$ of $s_{i}$ by Eve using $\mathbf{y}_{E}\doteq[\mathbf{y}_{E1}^{T},\cdots,\mathbf{y}_{EM}^{T},\mathbf{y}_{EA}^{T}]^{T}$ is
$\sigma_{\Delta s_{iE}}^{2}=1-\mathbf{r}_{i}^{H}\mathbf{R}^{-1}\mathbf{r}_{i}$ (90)
where $\mathbf{r}_{i}^{H}=\mathbb{E}\{s_{i}\mathbf{y}_{E}^{H}\}$ and $\mathbf{R}=\mathbb{E}\{\mathbf{y}_{E}\mathbf{y}_{E}^{H}\}$. Furthermore,
$\mathbf{r}_{i}=\left[\mathbf{0}_{n_{E}(i-1)}^{T},\sqrt{1/2}\mathbf{h}_{Ei}^{\prime T},\mathbf{0}_{n_{E}(M-i+1)}^{T}\right]^{T}$ (91)
and
$\mathbf{R}=\left[\begin{array}[]{ccc}\mathbf{R}_{1,1}&\cdots&\mathbf{R}_{1,M+1}\\ \cdots&\cdots&\cdots\\ \mathbf{R}_{M+1,1}&\cdots&\mathbf{R}_{M+1,M+1}\end{array}\right].$ (92)
Here $\mathbf{0}_{m}$ is a zero vector of $m$ elements, and $\mathbf{R}_{i,j}=\mathbf{R}_{j,i}^{H}$ for all $i$ and $j$. For $1\leq i\leq M$, $1\leq j\leq M$ and $i\neq j$,
$\mathbf{R}_{i,i}=(1/2)(1+c_{i})\mathbf{h}_{Ei}^{\prime}\mathbf{h}_{Ei}^{\prime H}+\mathbf{I}_{n_{E}},$ (93)
$\mathbf{R}_{i,j}=(1/2)\epsilon_{i,j}\mathbf{h}_{Ei}^{\prime}\mathbf{h}_{Ej}^{\prime H},$ (94)
$\mathbf{R}_{i,M+1}=\sqrt{1/2}\mathbf{h}_{Ei}^{\prime}\mathbf{r}_{x,i}^{H}\mathbf{H}_{EA}^{\prime H},$ (95)
$\mathbf{R}_{M+1,M+1}=\mathbf{H}_{EA}^{\prime}\mathbf{H}_{EA}^{\prime H}+\mathbf{I}_{n_{E}},$ (96)
where $\epsilon_{i,j}=\mathbb{E}\{\hat{p}_{i}\hat{p}_{j}^{*}\}=c_{i}c_{j}\phi_{i,j}$ and $\mathbf{r}_{x,i}=\mathbb{E}\{\mathbf{x}\hat{p}_{i}^{*}\}=c_{i}\mathbf{\bar{h}}_{i}^{*}$.
To obtain an insight into $\sigma_{\Delta s_{i,E}}^{2}$, let us next choose $i=1$ without loss of generality. We can rewrite (92) as
$\mathbf{R}=\left[\begin{array}[]{cc}\mathbf{R}_{1,1}&\mathbf{\bar{R}}_{1}\\ \mathbf{\bar{R}}_{1}^{H}&\mathbf{\bar{R}}_{1,1}\end{array}\right]$ (97)
where $\mathbf{R}_{1,1}$ is the same $n_{E}\times n_{E}$ upper-left block of $\mathbf{R}$ in (92). Then
$\mathbf{R}^{-1}=\left[\begin{array}[]{cc}(\mathbf{R}_{1,1}-\mathbf{\bar{R}}_{1}\mathbf{\bar{R}}_{1,1}^{-1}\mathbf{\bar{R}}_{1}^{H})^{-1}&*\\ &*\end{array}\right]$ (98)
where * denotes matrix blocks of no importance. Hence, (90) with $i=1$ becomes
$\sigma_{\Delta s_{1,E}}^{2}=1-(1/2)\mathbf{h}_{E1}^{\prime H}(\mathbf{R}_{1,1}-\mathbf{\bar{R}}_{1}\mathbf{\bar{R}}_{1,1}^{-1}\mathbf{\bar{R}}_{1}^{H})^{-1}\mathbf{h}_{E1}^{\prime}.$ (99)
Recall
$\mathbf{R}_{1,1}=(1/2)(1+c_{1})\mathbf{h}_{E1}^{\prime}\mathbf{h}_{E1}^{\prime H}+\mathbf{I}_{n_{E}},$ (100)
$\mathbf{\bar{R}}_{1}=\sqrt{1/2}\mathbf{h}_{E1}^{\prime}\mathbf{c}_{1}^{H},$ (101)
with $\mathbf{c}_{1}^{H}=\left[\sqrt{1/2}\epsilon_{1,2}\mathbf{h}_{E2}^{\prime H},\cdots,\sqrt{1/2}\epsilon_{1,M}\mathbf{h}_{EM}^{\prime H},\mathbf{r}_{x,1}^{H}\mathbf{H}_{EA}^{\prime H}\right]$. Hence
$\mathbf{\bar{R}}_{1}\mathbf{\bar{R}}_{1,1}^{-1}\mathbf{\bar{R}}_{1}^{H}=(1/2)\mathbf{h}_{E1}^{\prime}\mathbf{c}_{1}^{H}\mathbf{\bar{R}}_{1,1}^{-1}\mathbf{c}_{1}\mathbf{h}_{E1}^{\prime H}.$ (102)
Let
$\gamma_{1}=1+c_{1}-\mathbf{c}_{1}^{H}\mathbf{\bar{R}}_{1,1}^{-1}\mathbf{c}_{1}.$ (103)
We see that $\gamma_{1}-1=c_{1}-\mathbf{c}_{1}^{H}\mathbf{\bar{R}}_{1,1}^{-1}\mathbf{c}_{1}>0$ which is effectively the MSE of the MMSE estimate of $\hat{p}_{1}$ by Eve using $\mathbf{y}_{E|1}\doteq[\mathbf{y}_{E,2}^{T},\cdots,\mathbf{y}_{E,M}^{T},\mathbf{y}_{EA}^{T}]^{T}$.
It follows from (99) that
$\sigma_{\Delta s_{1,E}}^{2}=1-(1/2)\mathbf{h}_{E1}^{\prime H}\left((1/2)\gamma_{1}\mathbf{h}_{E1}^{\prime}\mathbf{h}_{E1}^{\prime H}+\mathbf{I}_{n_{E}}\right)^{-1}\mathbf{h}_{E1}^{\prime}$
$=1-\left(\gamma_{1}\|\mathbf{h}_{E1}^{\prime}\|^{2}/2+1\right)^{-1}\|\mathbf{h}_{E1}^{\prime}\|^{2}/2$
$=\frac{(\gamma_{1}-1)S_{E,1}/2+1}{\gamma_{1}S_{E,1}/2+1},$ (104)
with $S_{E,1}=\|\mathbf{h}_{E1}^{\prime}\|^{2}$.
Finally, the capacity of the effective return channel from an arbitrary UE, labeled as UE1, to AP relative to $s_{1}$ is
$C_{E|1}\doteq I(s_{1};\mathbf{y}_{E})=I(s_{1};\hat{s}_{1,E})=\log(1/\sigma_{\Delta s_{1,E}}^{2})$
$=\log\left(1+\frac{S_{E,1}/2}{(\gamma_{1}-1)S_{E,1}/2+1}\right).$ (105)
Proposition 4
An achievable secrecy rate of M-STEEP from an arbitrarily selected UE1 to the AP relative to the message symbol $s_{1}$ from UE1 is
$C_{s,1}=(C_{A|1}-C_{E|1})^{+}=\left[\log\left(1+\frac{\frac{S_{A,1}}{2}}{\frac{S_{1}S_{A,1}/2}{(S_{1}+1)^{2}}+1}\right)\right.$
$\,\,\left.-\log\left(1+\frac{S_{E,1}/2}{(\gamma_{1}-1)S_{E,1}/2+1}\right)\right]^{+}$ (106)
where only $\gamma_{1}$ is affected by UEi’s power for all $i$, i.e., only $\gamma_{1}$ depends on $S_{E,i}=\|\mathbf{h}_{Ei}^{\prime}\|^{2}$ for all $i=1,\cdots,M$.
This proposition follows from (87) and (V-A). If $M=1$, it reduces to Corollary 1.
Proposition 5
Assume $n_{A}=1$, rewrite $\mathbf{H}_{EA}^{\prime}$ as $\mathbf{h}_{EA}^{\prime}$, and let $S_{i}=\|\mathbf{h}_{i}^{\prime}\|^{2}$, $S_{Ai}=\|\mathbf{h}_{Ai}^{\prime}\|^{2}$ and $S_{Ei}=\|\mathbf{h}_{Ei}^{\prime}\|^{2}$, $S_{EA}=\|\mathbf{h}_{EA}^{\prime}\|^{2}$, $\alpha_{i}=\frac{S_{EA}}{S_{i}}$ and $\beta_{i}=\frac{S_{Ei}}{S_{Ai}}$. Then $\gamma_{1}-1$ in (4) becomes
$\gamma_{1}-1=\frac{S_{1}}{(S_{1}+1)^{2}}\left(1+\frac{S_{1}}{\alpha_{1}S_{1}+1}\left(1-\frac{t_{1,M}}{\alpha_{1}S_{1}+1}\right)\right)$ (107)
Also $t_{1,M}=0$ for $M=1$, and $t_{1,M}$ for $M\geq 2$ is defined in (162) which is a function of $S_{E,A}$ and $S_{E,i}$ for all $i\neq 1$. And $t_{1,M}<\min(M-1,\alpha_{1}S_{1}+1)$. Consequently, $C_{s,1}>0$ if and only if
$S_{A,1}/2>\left(1-\frac{1}{\beta_{1}}\right)\frac{(S_{1}+1)^{2}(\alpha_{1}S_{1}+1)}{S_{1}^{2}\left(1-\frac{t_{1,M}}{\alpha_{1}S_{1}+1}\right)}.$ (108)
Note that the left side of (108) is proportional to $p_{u1}$ (the power from UE1) and the right side of (108) is invariant to $p_{u1}$ and large $p_{ui}$ for all $i\neq 1$.
Proof:
See Appendix-C. ∎
This proposition has also been validated by computer simulations. If $M=1$, (108) reduces to (3). But more importantly, we see from (108) that for any given $M$, the secrecy rate from any UE to AP stays positive if that UE uses a sufficiently large power according to (108).
V-B Total Secrecy Capacity of M-STEEP
A total secrecy rate of M-STEEP can be expressed as
$C_{s}\doteq C_{s,1}+C_{s,2|1}+\cdots+C_{s,M|1,\cdots,M-1}$ (109)
with $C_{s,i|1,\cdots,i-1}\doteq[C_{A,i|1,\cdots,i-1}-C_{E,i|1,\cdots,i-1}]^{+}$ which is the secrecy rate from EUi to AP subject to $s_{1},\cdots,s_{i-1}$ being known to AP and Eve. Furthermore, we have $C_{A,i|1,\cdots,i-1}=I(s_{i};\mathbf{y}_{A}|s_{1},\cdots,s_{i-1},\mathbf{x})$ and $C_{E,i|1,\cdots,i-1}=I(s_{i};\mathbf{y}_{E}|s_{1},\cdots,s_{i-1})$. Here $\mathbf{y}_{A}=[\mathbf{y}_{A1}^{T},\cdots,\mathbf{y}_{AM}^{T}]^{T}$ and $\mathbf{y}_{E}=[\mathbf{y}_{E1}^{T},\cdots,\mathbf{y}_{EM}^{T},\mathbf{y}_{EA}^{T}]^{T}$.
It is easy to verify that the capacity from UEi to AP for $s_{i}$ is invariant to $s_{1},\cdots,s_{i-1}$ being known to AP or not. Hence $C_{A,i|1,\cdots,i-1}=C_{A,i}$ as given by (87).
The MSE of the MMSE estimate of $s_{i}$ by Eve using $\mathbf{y}_{E}$ and $s_{1},\cdots,s_{i-1}$ is
$\sigma_{\Delta s_{i,E|1,\cdots,i-1}}^{2}=1-\mathbf{r}_{i}^{H}\mathbf{R}_{i}^{-1}\mathbf{r}_{i}$ (110)
where $\mathbf{r}_{i}^{H}=\mathbb{E}\{s_{i}\mathbf{y}_{E|1,\cdots,i-1}^{H}\}$ and $\mathbf{R}_{i}=\mathbb{E}\{\mathbf{y}_{E|1,\cdots,i-1}\mathbf{y}_{E|1,\cdots,i-1}^{H}\}$. Here
$\mathbf{y}_{E|1,\cdots,i-1}\doteq$
$\,\,[\mathbf{\bar{y}}_{E,1}^{T},\cdots,\mathbf{\bar{y}}_{E,i-1}^{T},\mathbf{y}_{E,i}^{T},\cdots,\mathbf{y}_{E,M}^{T},\mathbf{y}_{E,A}^{T}]^{T}$ (111)
with $\mathbf{\bar{y}}_{E,j}=\mathbf{y}_{E,j}-\mathbf{h}_{Ej}^{\prime}s_{j}$. Furthermore, $\mathbf{r}_{i}$ is given by (91), and $\mathbf{R}_{i}$ is the same as $\mathbf{R}$ in (92) except that the $j$th diagonal block $\mathbf{R}_{j,j}$ of $\mathbf{R}$ for $j=1,\cdots,i-1$ should be replaced by $\mathbf{\bar{R}}_{j,j}=c_{j}\mathbf{h}_{Ej}^{\prime}\mathbf{h}_{Ej}^{\prime H}+\mathbf{I}_{n_{E}}$. Finally, $C_{E,i|1,\cdots,i-1}=-\log\sigma_{\Delta s_{i,E|1,\cdots,i-1}}^{2}$.
Subject to i.i.d. Gaussian random channel parameters, numerical results [30] suggest that the medium of $C_{s}$ increases as $M$ increases while the medium of $C_{s,1}$ decreases (but rather slowly) as $M$ increases. Further mathematical analysis of $C_{s}$ remains open.
VI Additional Comments
An important connection between STEEP and the works in [7], [2], [28] and [22] has been stressed in the introduction. The idea of feedback for WTC shown in [27] however has a limited applicability unlike STEEP.
The channel probing idea used for phase 1 of STEEP was initially inspired from [13] and [14] where the authors attempted to use random channel probing to increase the secret-key rate from static reciprocal channels (but with no proven success). This idea is indeed only a special case of the notion of generating correlated data sets at Alice and Bob for SKG [7]. The correlated data sets can be directly generated from a channel model as shown in Chapter 4 in [2]. However, the treatment in [2] is mostly on the validity and meaning of secret-key capacity and its bounds. It does not address the application of these bounds. The works shown in [23] and [24] however addressed SKG in more applied settings using data sets collected from some specific channel models as well as public communications.
To achieve the secrecy capacity of STEEP, WTC coding is needed in the echoing phase (phase 2). Such coding schemes are available, including LDPC codes [19] and polar codes [18]. However, the current practicality of those codes due to their complexity seems questionable.
But the effective WTC system constructed by STEEP is such that the user’s effective channel is almost surely stronger than Eve’s effective channel. Because of this, a positive secrecy rate is virtually given without the need to know Eve’s CSI. Furthermore, to realize a positive secrecy rate for STEEP, we do not necessarily need to use a capacity-achieving channel coding scheme. All we need is a channel code for which the optimal decoding can be done by the receiving user in phase 2. For example, a convolution code can be used by Bob in phase 2 to encode the stream of the secret information (e.g., $\mathbf{s}$ in (9) or $\phi$ in (59)). Then Alice can perform the maximum likelihood decoding, such as Viterbi decoding, of the secret information (e.g., from $\mathbf{\hat{s}}$ in (11) or $r_{A}$ in (60)). Since the decoding at Alice is optimal and the effective channel from Bob to Alice is stronger than the effective channel from Bob to Eve, the error rate at Eve is always higher than that at Alice. The lack of capacity achieving of a channel code would reduce the net channel capacities for both user and Eve but without necessarily a significant change to a positive secrecy rate. For a Gaussian-noise channel, the error rate drops exponentially as SNR increases, which creates a drastic difference between the number of errors at Alice and those at Eve. Such a gap of error rates can be used as a secrecy measure.
Provided no error is detected at Alice (using any of the established channel codes), if the secret information is meant to generate a secret key, a hash function could then be applied at Alice and Bob to produce the secret key with a higher confidence of its secrecy (also known as privacy amplification). The secret-key rate in bits/s/Hz of this STEEP-assisted method for SKG does not reduce to zero as the channel coherence time increases, unlike numerous methods in the literature such as [20] and [21] based on reciprocal channels. To know the exact amount of secrecy, it would always require the knowledge of Eve’s channel. But it could suffice in practice that there is at least some amount of positive secrecy rate even in the worst possible case.
STEEP may also remind one of a widely used method for networking security called “nonce”. The usefulness of nonce is based on the assumption that Alice can send a nonce reliably to Bob while Eve can not receive it. Then this nonce can be used (normally once) by Bob to encrypt a message to be sent to Alice. Unlike nonce, STEEP allows Eve to receive the probes from Alice but with some noise while Bob does not have to receive the probes with more accuracy than Eve, and the noisy probes received by Bob are used to encrypt a secret message to be sent to Alice. STEEP is naturally applicable at the physical layer due to presence of independent noises while it is also applicable at a higher layer.
VII Conclusion
This paper has presented STEEP in its latest forms, and demonstrated a broad applicability of STEEP for secure communications especially at the physical layer. Its connection to some prior works such as in [7], [2], [28], [22] has been highlighted. Novel insights into how STEEP can be further formulated in different settings have been revealed. Furthermore, the achievable secrecy rates of STEEP in three different settings of practical importance have been derived and analyzed. In each of the settings, STEEP is shown to yield a positive secrecy rate, subject to a sufficiently large power in the echoing phase, as long as Eve’s receive channel in the probing phase is not noiseless. This attractive property of STEEP has its roots in the encryption lemma discussed in the introduction. Furthermore, the broad applicability of STEEP should make it a strong candidate for secure communications in practice - towards a full realization of a hope in [7].
Appendix
-A Proof of Proposition 1
We will assume $n_{A}\geq n_{B}$. Recall $\mathbf{R}_{\Delta\mathbf{p}^{\prime}}$ in (III-B), $\mathbf{R}_{\Delta\mathbf{p}}$ in (III-A) and $C_{A|B,G}=\log\frac{N_{A|B}}{D_{A|B}}$ in (III-C). Also recall $\mathbf{H}_{AB}^{\prime\prime}=\sqrt{p_{B}/(2n_{B})}\mathbf{H}_{AB}$, $\mathbf{H}_{EB}^{\prime\prime}=\sqrt{p_{B}/(2n_{B})}\mathbf{H}_{EB}$, $\mathbf{H}_{BA}^{\prime}=\sqrt{p_{A}/n_{A}}\mathbf{H}_{BA}$, and $\mathbf{H}_{EA}^{\prime}=\sqrt{p_{A}/n_{A}}\mathbf{H}_{EA}$.
We know $\lim_{p_{A}\to\infty}\mathbf{R}_{\Delta\mathbf{p}^{\prime}}=\lim_{p_{A}\to\infty}\mathbf{R}_{\Delta\mathbf{p}}=\mathbf{0}$. Then, subject to a fixed $\eta_{p}=\frac{p_{B}}{p_{A}}$,
$\lim_{p_{A}\to\infty}\log D_{A|B}$
$=\log\left|\frac{\eta_{p}n_{A}}{2n_{B}}\mathbf{H}_{AB}^{H}\mathbf{H}_{AB}(\boldsymbol{\Pi}_{BA}^{2})^{-1}+\mathbf{I}_{n_{B}}\right|$ (112)
which is invariant to $p_{A}$, and
$\lim_{p_{A}\to\infty}\log N_{A|B}$
$=\lim_{p_{A}\to\infty}\log\left|\frac{p_{B}}{2n_{B}}\mathbf{H}_{AB}^{H}\mathbf{H}_{AB}\right|+o(\log p_{A}).$ (113)
Here $o(x)$ is a quantity such that $\lim_{x\to\infty}\frac{o(x)}{x}=0$. And therefore, for $n_{A}\geq n_{B}$ and a fixed $\eta_{p}=\frac{p_{B}}{p_{A}}$,
$\lim_{p_{A}\to\infty}\log N_{A|B}/\log p_{A}=n_{B},$ (114)
$\lim_{p_{A}\to\infty}\log D_{A|B}/\log p_{A}=0,$ (115)
and hence
$\lim_{p_{A}\to\infty}C_{A|B,G}/\log p_{A}=n_{B}.$ (116)
Now let us consider $C_{E|B,G}=\log\frac{N_{E|B}}{D_{E|B}}$ in (III-D) and $\mathbf{T}$ in (III-D). We know that $rank(\mathbf{H}_{EA})=\min(n_{A},n_{E})\doteq r_{A}$. We can write the eigenvalue decomposition (EVD) of $\mathbf{H}_{EA}^{H}\mathbf{H}_{EA}$ as $\mathbf{U}\mathbf{D}^{2}\mathbf{U}^{H}=\mathbf{U}_{A}\mathbf{D}_{A}^{2}\mathbf{U}_{A}^{H}$ where $\mathbf{D}_{A}^{2}$ is $r_{A}\times r_{A}$ nonsingular diagonal and $\mathbf{U}_{A}$ is the corresponding $r_{A}$ columns of the $n_{A}\times n_{A}$ unitary matrix $\mathbf{U}$. It follows from (III-D) that
$\mathbf{T}=\frac{p_{A}}{n_{A}}\mathbf{R}_{\mathbf{\hat{p}}}^{H}\mathbf{V}_{BA}^{H}\mathbf{U}\left(\frac{p_{A}}{n_{A}}\mathbf{D}^{2}+\mathbf{I}_{n_{A}}\right)^{-1}$
$\,\,\cdot\mathbf{D}^{2}\mathbf{U}^{H}\mathbf{V}_{BA}\mathbf{R}_{\mathbf{\hat{p}}}$
$=\frac{p_{A}}{n_{A}}\mathbf{R}_{\mathbf{\hat{p}}}^{H}\mathbf{V}_{BA}^{H}\mathbf{U}_{A}\left(\frac{p_{A}}{n_{A}}\mathbf{D}_{A}^{2}+\mathbf{I}_{r_{A}}\right)^{-1}$
$\,\,\cdot\mathbf{D}_{A}^{2}\mathbf{U}_{A}^{H}\mathbf{V}_{BA}\mathbf{R}_{\mathbf{\hat{p}}}.$ (117)
Hence,
$\lim_{p_{A}\to\infty}\mathbf{T}=\mathbf{V}_{BA}^{H}\mathbf{U}_{A}\mathbf{U}_{A}^{H}\mathbf{V}_{BA}$ (118)
where we have used $\lim_{p_{A}\to\infty}\mathbf{R}_{\mathbf{\hat{p}}}=\mathbf{I}_{n_{B}}$.
Since $\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}=\mathbf{R}_{\mathbf{\hat{p}}}-\mathbf{T}$ as in (40), then for any $p_{B}$,
$\lim_{p_{A}\to\infty}\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}=\mathbf{I}_{n_{B}}-\mathbf{V}_{BA}^{H}\mathbf{U}_{A}\mathbf{U}_{A}^{H}\mathbf{V}_{BA}$
$=\mathbf{V}_{BA}^{H}\mathbf{P}_{A}^{\perp}\mathbf{V}_{BA},$ (119)
where $\mathbf{P}_{A}^{\perp}=\mathbf{I}_{n_{A}}-\mathbf{U}_{A}\mathbf{U}_{A}^{H}$ is the projection matrix onto the orthogonal complement of $range(\mathbf{U}_{A})$, and has the rank $(n_{A}-n_{E})^{+}$. Furthermore,
$\lim_{p_{A}\to\infty}\log D_{E|B}$
$=\log\left|\frac{p_{B}}{2n_{B}}\mathbf{H}_{EB}^{H}\mathbf{H}_{EB}\mathbf{V}_{BA}^{H}\mathbf{P}_{A}^{\perp}\mathbf{V}_{BA}+\mathbf{I}_{n_{B}}\right|,$ (120)
$\lim_{p_{A}\to\infty}\log N_{E|B}$
$=\log\left|\frac{p_{B}}{2n_{B}}\mathbf{H}_{EB}^{H}\mathbf{H}_{EB}(\mathbf{V}_{BA}^{H}\mathbf{P}_{A}^{\perp}\mathbf{V}_{BA}+\mathbf{I}_{n_{B}})+\mathbf{I}_{n_{B}}\right|.$ (121)
It is typical (or with probability one for random matrices) that
$rank(\mathbf{H}_{EB}^{H}\mathbf{H}_{EB}\mathbf{V}_{BA}^{H}\mathbf{P}_{A}^{\perp}\mathbf{V}_{BA})$
$=\min(rank(\mathbf{H}_{EB}),rank(\mathbf{V}_{BA}),rank(\mathbf{P}_{A}^{\perp}))$
$=\min(n_{B},(n_{A}-n_{E})^{+}),$ (122)
and
$rank(\mathbf{H}_{EB}^{H}\mathbf{H}_{EB}\mathbf{V}_{BA}^{H}\mathbf{P}_{A}^{\perp}\mathbf{V}_{BA}+\mathbf{I}_{n_{B}})=n_{B}.$ (123)
Therefore, for fixed $\eta_{p}=\frac{p_{B}}{p_{A}}$,
$\lim_{p_{A}\to\infty}N_{E|B}/\log p_{A}=n_{B},$ (124)
$\lim_{p_{A}\to\infty}D_{E|B}/\log p_{A}=\min(n_{B},(n_{A}-n_{E})^{+}),$ (125)
and hence
$\lim_{p_{A}\to\infty}\frac{C_{E|B,G}}{\log p_{A}}=n_{B}-\min(n_{B},(n_{A}-n_{E})^{+}).$ (126)
Applying (116) and (126) yields that for fixed $\eta_{p}=\frac{p_{B}}{p_{A}}$,
$\lim_{p_{A}\to\infty}\frac{1}{\log p_{A}}C_{s,G}=\min(n_{B},(n_{A}-n_{E})^{+}).$ (127)
Finally, we rewrite (47) as
$C_{key}=\log\left|\mathbf{I}_{n_{A}}+\frac{p_{A}}{n_{A}}\mathbf{\tilde{H}}_{EA}^{H}\mathbf{\tilde{H}}_{EA}\right|$
$\,\,-\log\left|\frac{p_{A}}{n_{A}}\mathbf{H}_{EA}^{H}\mathbf{H}_{EA}+\mathbf{I}_{n_{A}}\right|$ (128)
with $\mathbf{\tilde{H}}_{EA}=\left[\begin{array}[]{c}\mathbf{H}_{EA}\\ \mathbf{H}_{BA}\end{array}\right]$. Here, $rank(\mathbf{H}_{EA})=\min(n_{A},n_{E})$ and
$rank(\mathbf{\tilde{H}}_{EA})=\min(n_{A},n_{E}+n_{B}).$ (129)
It follows that
$\lim_{p_{A}\to\infty}\frac{1}{\log p_{A}}C_{key}=\min(n_{A},n_{E}+n_{B})-\min(n_{A},n_{E})$
$=\min(n_{B},(n_{A}-n_{E})^{+}).$ (130)
The proof is completed.
-B Proof of Proposition 2
It is easy to verify that $C_{key}$ from (47) satisfies the second equation in (2). We will next show the first equation in (2).
For $n_{E}\geq n_{A}\geq n_{B}$, both $\mathbf{H}_{AB}^{H}\mathbf{H}_{AB}$ and $\mathbf{H}_{EB}^{H}\mathbf{H}_{EB}$ have full rank $n_{B}$, and hence (1) implies
$C_{s,G}^{B}\doteq\lim_{p_{B}\to\infty}C_{s,G}$
$=\log\left(\frac{|(\mathbf{R}_{\Delta\mathbf{p}^{\prime}}+\mathbf{I}_{n_{B}})|}{|\mathbf{R}_{\Delta\mathbf{p}^{\prime}}|}\frac{|\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}|}{|(\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}+\mathbf{I}_{n_{B}})|}\right).$ (131)
Next we consider $C_{s,G}^{B,A}\doteq\lim_{p_{A}\to\infty}C_{s,G}^{B}$. Since $\mathbf{H}_{EA}^{H}\mathbf{H}_{EA}$ is invertible, then $\lim_{p_{A}\to\infty}\mathbf{T}=\mathbf{I}_{n_{B}}$ and $\lim_{p_{A}\to\infty}\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}=0$. Also note that $\mathbf{R}_{\mathbf{\hat{p}}}=\mathbf{I}_{n_{B}}+\mathcal{O}(1/p_{A})$ and $\lim_{p_{A}\to\infty}\mathbf{R}_{\Delta\mathbf{p}^{\prime}}=\lim_{p_{A}\to\infty}\mathbf{R}_{\Delta\mathbf{p}}=0$. Therefore,
$C_{s,G}^{B,A}\doteq\lim_{p_{A}\to\infty}C_{s,G}^{B}$
$=\lim_{p_{A}\to\infty}\log\left(\frac{|\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}|}{|\mathbf{R}_{\Delta\mathbf{p}}|}\right)$ (132)
where $\mathbf{R}_{\Delta\mathbf{p}^{\prime}}=\mathbf{R}_{\Delta\mathbf{p}}(\mathbf{I}+\mathcal{O}(1/p_{A}))$ for large $p_{A}$ has been used, and the indefinite form of $\frac{0}{0}$ is resolved next.
Let $\gamma_{A}\doteq\frac{p_{A}}{n_{A}}$, $\mathbf{W}_{E}\doteq\mathbf{H}_{EA}^{H}\mathbf{H}_{EA}$ and $\mathbf{T}_{P}\doteq\gamma_{A}\boldsymbol{\Pi}_{BA}^{2}+\mathbf{I}_{n_{B}}$. Then
$\mathbf{R}_{\Delta\mathbf{p}}^{-1}\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}=\mathbf{T}_{P}\left(\gamma_{A}\boldsymbol{\Pi}_{BA}^{2}\mathbf{T}_{P}^{-1}\right.$
$\,\,\left.-\gamma_{A}\mathbf{R}_{\mathbf{\hat{p}}}^{H}\mathbf{V}_{BA}^{H}\left(\gamma_{A}\mathbf{W}_{E}+\mathbf{I}_{n_{E}}\right)^{-1}\mathbf{W}_{E}\mathbf{V}_{BA}\mathbf{R}_{\mathbf{\hat{p}}}\right)$
$=\gamma_{A}\boldsymbol{\Pi}_{BA}^{2}-\gamma_{A}^{2}\boldsymbol{\Pi}_{BA}^{2}\mathbf{V}_{BA}^{H}\left(\gamma_{A}\mathbf{W}_{E}+\mathbf{I}_{n_{E}}\right)^{-1}$
$\,\,\cdot\mathbf{W}_{E}\mathbf{V}_{BA}\mathbf{R}_{\mathbf{\hat{p}}}$ (133)
Also note
$(\gamma_{A}\mathbf{W}_{E}+\mathbf{I}_{n_{E}})^{-1}=\gamma_{A}^{-1}\mathbf{W}_{E}^{-1}$
$\,\,-\gamma_{A}^{-1}\mathbf{W}_{E}^{-1}(\mathbf{I}_{n_{E}}+\gamma_{A}^{-1}\mathbf{W}_{E}^{-1})^{-1}\gamma_{A}^{-1}\mathbf{W}_{E}^{-1},$ (134)
$\mathbf{R}_{\mathbf{\hat{p}}}=\gamma_{A}\boldsymbol{\Pi}_{BA}^{2}(\gamma_{A}\boldsymbol{\Pi}_{BA}^{2}+\mathbf{I}_{n_{B}})^{-1}$
$=\mathbf{I}_{n_{B}}-(\mathbf{I}_{n_{B}}+\gamma_{A}^{-1}\boldsymbol{\Pi}_{BA}^{-2})^{-1}\gamma_{A}^{-1}\boldsymbol{\Pi}_{BA}^{-2}.$ (135)
Then, one can verify that
$\mathbf{R}_{\Delta\mathbf{p}}^{-1}\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}=(\mathbf{I}_{n_{B}}+\gamma_{A}^{-1}\boldsymbol{\Pi}_{BA}^{-2})^{-1}+$
$\,\,\boldsymbol{\Pi}_{BA}^{2}\mathbf{V}_{BA}^{H}\mathbf{W}_{E}^{-1}(\mathbf{I}_{n_{A}}+\gamma_{A}^{-1}\mathbf{W}_{E}^{-1})^{-1}\mathbf{V}_{BA}\mathbf{R}_{\mathbf{\hat{p}}},$ (136)
and therefore
$C_{s,G}^{B,A}=\lim_{p_{A}\to\infty}\log|\mathbf{R}_{\Delta\mathbf{p}}^{-1}\mathbf{R}_{\Delta\mathbf{\hat{p}}_{E}}|$
$=\log\left|\mathbf{I}_{n_{B}}+\boldsymbol{\Pi}_{BA}^{2}\mathbf{V}_{BA}^{H}\mathbf{W}_{E}^{-1}\mathbf{V}_{BA}\right|$ (137)
which completes the proof.
-C Proof of Proposition 5
It is easy to verify that subject to (107), $C_{s,1}>0$ if and only if
$\frac{S_{A,1}}{2}\left(1-\frac{t_{1,M}}{\alpha_{1}S_{1}+1}\right)>\left(1-\frac{1}{\beta_{1}}\right)\frac{(S_{1}+1)^{2}(\alpha_{1}S_{1}+1)}{S_{1}^{2}}.$ (138)
It will be obvious that for $M=1$, $t_{1,M}=0$. We will also see that $t_{1,M}$ is an increasing function of $S_{E,A}$ while $\frac{t_{1,M}}{S_{E,A}+1}$ is a decreasing function of $S_{E,A}$.
We will also show explicitly that $t_{1,M}<M-1$. This means that if $S_{E,A}=\alpha_{1}S_{1}>M-2$, then $01-\frac{t_{1,M}}{\alpha_{1}S_{1}+1}>$ 0 and hence (138) is equivalent to (108). Since $C_{s,1}$ must be a decreasing function of $\alpha_{1}$ (which is the ratio of the channel strength from AP to Eve over that from AP to UE1), $C_{s,1}$ must increase as $\alpha_{1}$ decreases. If the peak value of $\frac{t_{1,M}}{\alpha_{1}S_{1}+1}$ (a decreasing function of $\alpha_{1}$) is larger than one, then as $\alpha_{1}$ decreases (starting from the condition $S_{E,A}=\alpha_{1}S_{1}>M-2$) the condition $01-\frac{t_{1,M}}{\alpha_{1}S_{1}+1}>$ 0 would be reversed and a contradiction from $C_{s,1}>0$ would be concluded. Therefore, $t_{1,M}<\alpha_{1}S_{1}+1$ must hold for all $\alpha_{1}$. This also means that $t_{1,M}<1$ when $S_{E,A}=\alpha_{1}S_{1}=0$.
We will only need to prove (107) for all $M$, and $t_{1,M}<M-1$ for $M\geq 2$. Recall $\gamma_{1}-1=c_{1}-q_{1}$ with
$q_{1}=\mathbf{c}_{1}^{H}\mathbf{\bar{R}}_{1,1}^{-1}\mathbf{c}_{1}.$ (139)
Also recall $c_{i}=\sigma_{\hat{p}_{i}}^{2}=\mathcal{E}\{|\hat{p}_{i}|^{2}\}$; and for $n_{A}=1$, $\phi_{i,j}=1$ and hence $\epsilon_{i,j}=c_{i}c_{j}$.
To simplify the notions, we will use $\mathbf{g}_{i}\doteq\sqrt{1/2}\mathbf{h}_{Ei}^{\prime}$ and $\mathbf{g}_{A}\doteq\mathbf{h}_{EA}^{\prime}$. Then,
$\mathbf{c}_{1}^{H}=c_{1}[c_{2}\mathbf{g}_{2}^{H},c_{3}\mathbf{g}_{3}^{H},\cdots,c_{M}\mathbf{g}_{M}^{H}|\mathbf{g}_{A}^{H}]\doteq c_{1}[\mathbf{c}_{a}^{H}|\mathbf{c}_{b}^{H}]$ (140)
with $\mathbf{c}_{a}=[c_{2}\mathbf{g}_{2}^{H},c_{3}\mathbf{g}_{3}^{H},\cdots,c_{M}\mathbf{g}_{M}^{H}]^{H}$ and $\mathbf{c}_{b}=\mathbf{g}_{A}$. Also
$\mathbf{\bar{R}}_{1,1}=\mathbf{I}+$
$\left[\begin{array}[]{ccc|c}(1+c_{2}^{2})\mathbf{g}_{2}\mathbf{g}_{2}^{H}&\cdots&c_{2}c_{M}\mathbf{g}_{2}\mathbf{g}_{M}^{H}&c_{2}\mathbf{g}_{2}\mathbf{g}_{A}^{H}\\ \cdots&\cdots&\cdots&\cdots\\ c_{M}c_{2}\mathbf{g}_{M}\mathbf{g}_{2}^{H}&\cdots&(1+c_{M}^{2})\mathbf{g}_{M}\mathbf{g}_{M}^{H}&c_{M}\mathbf{g}_{M}\mathbf{g}_{A}^{H}\\ \hline\cr c_{2}\mathbf{g}_{A}\mathbf{g}_{2}^{H}&\cdots&c_{M}\mathbf{g}_{A}\mathbf{g}_{M}^{H}&\mathbf{g}_{A}\mathbf{g}_{A}^{H}\end{array}\right]$ (145)
$\doteq\left[\begin{array}[]{c|c}\mathbf{A}_{1,1}&\mathbf{A}_{1,2}\\ \hline\cr\mathbf{A}_{1,2}^{H}&\mathbf{A}_{2,2}\end{array}\right]$ (148)
with $\mathbf{A}_{1,2}=\mathbf{c}_{a}\mathbf{g}_{A}^{H}$ and $\mathbf{A}_{2,2}=\mathbf{g}_{A}\mathbf{g}_{A}^{H}+\mathbf{I}$. Also let
$\mathbf{\bar{R}}_{1,1}^{-1}=\left[\begin{array}[]{cc}\mathbf{B}_{1,1}&\mathbf{B}_{1,2}\\ \mathbf{B}_{1,2}^{H}&\mathbf{B}_{2,2}\end{array}\right]$ (149)
Then it is known that
$\mathbf{B}_{1,1}=(\mathbf{A}_{1,1}-\mathbf{A}_{1,2}\mathbf{A}_{2,2}^{-1}\mathbf{A}_{1,2}^{H})^{-1},$ (150)
$\mathbf{B}_{2,2}=\mathbf{A}_{2,2}^{-1}+\mathbf{A}_{2,2}^{-1}\mathbf{A}_{1,2}^{H}\mathbf{B}_{1,1}\mathbf{A}_{1,2}\mathbf{A}_{2,2}^{-1},$ (151)
$\mathbf{B}_{1,2}=-\mathbf{B}_{1,1}\mathbf{A}_{1,2}\mathbf{A}_{2,2}^{-1}.$ (152)
Here
$\mathbf{A}_{1,2}\mathbf{A}_{2,2}^{-1}\mathbf{A}_{1,2}^{H}=\mathbf{c}_{a}\mathbf{g}_{A}^{H}(\mathbf{g}_{A}\mathbf{g}_{A}^{H}+\mathbf{I})^{-1}\mathbf{g}_{A}\mathbf{c}_{a}^{H}$
$=\frac{S_{E,A}}{S_{E,A}+1}\mathbf{c}_{a}\mathbf{c}_{a}^{H}$ (153)
with $S_{E,A}=\|\mathbf{g}_{A}\|^{2}$. Then, it follows from (150) that
$\mathbf{B}_{1,1}=\left(\mathbf{I}+\right.$
$\left.\left[\begin{array}[]{ccc}\left(1+\frac{c_{2}^{2}}{S_{E,A}+1}\right)\mathbf{g}_{2}\mathbf{g}_{2}^{H}&\cdots&\frac{c_{2}c_{M}}{S_{E,A}+1}\mathbf{g}_{2}\mathbf{g}_{M}^{H}\\ \cdots&\cdots&\cdots\\ \frac{c_{M}c_{2}}{S_{E,A}+1}\mathbf{g}_{M}\mathbf{g}_{2}^{H}&\cdots&\left(1+\frac{c_{M}^{2}}{S_{E,A}+1}\right)\mathbf{g}_{M}\mathbf{g}_{M}^{H}\end{array}\right]\right)^{-1}$ (157)
We know from (139), (140) and (149) that
$q_{1}=c_{1}^{2}(q_{a}+2\Re\{q_{b}\}+q_{c})$ (158)
with $q_{a}=\mathbf{c}_{a}^{H}\mathbf{B}_{1,1}\mathbf{c}_{a}$, $q_{b}=\mathbf{c}_{a}^{H}\mathbf{B}_{1,2}\mathbf{c}_{b}$ and $q_{c}=\mathbf{c}_{b}^{H}\mathbf{B}_{2,2}\mathbf{c}_{b}$.
It follows that
$q_{b}=-\mathbf{c}_{a}^{H}\mathbf{B}_{1,1}\mathbf{c}_{a}\mathbf{g}_{A}^{H}(\mathbf{g}_{A}\mathbf{g}_{A}^{H}+\mathbf{I})^{-1}\mathbf{g}_{A}$
$=-q_{a}\frac{S_{E,A}}{S_{E,A}+1}.$ (159)
And
$q_{c}=\mathbf{g}_{A}^{H}\left((\mathbf{g}_{A}\mathbf{g}_{A}^{H}+\mathbf{I})^{-1}+\right.$
$\,\,\left.(\mathbf{g}_{A}\mathbf{g}_{A}^{H}+\mathbf{I})^{-1}\mathbf{g}_{A}\mathbf{c}_{a}^{H}\mathbf{B}_{1,1}\mathbf{c}_{a}\mathbf{g}_{A}^{H}(\mathbf{g}_{A}\mathbf{g}_{A}^{H}+\mathbf{I})^{-1}\right)\mathbf{g}_{A}$
$=\left(\frac{S_{E,A}}{S_{E,A}+1}+\frac{S_{E,A}^{2}}{(S_{E,A}+1)^{2}}\mathbf{c}_{a}^{H}\mathbf{B}_{1,1}\mathbf{c}_{a}\right)$
$=\frac{S_{E,A}}{S_{E,A}+1}+\frac{S_{E,A}^{2}}{(S_{E,A}+1)^{2}}q_{a}.$ (160)
Therefore, (158) becomes
$q_{1}=c_{1}^{2}\left(q_{a}-2q_{a}\frac{S_{E,A}}{S_{E,A}+1}+\frac{S_{E,A}}{S_{E,A}+1}+\frac{S_{E,A}^{2}}{(S_{E,A}+1)^{2}}q_{a}\right)$
$=c_{1}^{2}\left(\frac{q_{a}}{(S_{E,A}+1)^{2}}+\frac{S_{E,A}}{S_{E,A}+1}\right).$ (161)
We now let $t_{1,M}\doteq q_{a}$. It is obvious that $t_{1,M}=0$ for $M=1$. We will show $t_{1,M}<M-1$ for $M\geq 2$. We can also write
$t_{1,M}=\mathbf{v}_{M}^{H}\mathbf{B}_{M}^{-1}\mathbf{v}_{M}$ (162)
with
$\mathbf{v}_{M}^{H}=[c_{2}\mathbf{g}_{2}^{H},\cdots,c_{M-1}\mathbf{g}_{M-1}^{H}|c_{M}\mathbf{g}_{M}^{H}]=[\mathbf{v}_{M-1}^{H}|c_{M}\mathbf{g}_{M}^{H}],$ (163)
$\mathbf{B}_{M}=\mathbf{I}+$
$\left[\begin{array}[]{cc|c}\left(1+\frac{c_{2}^{2}}{S_{E,A}+1}\right)\mathbf{g}_{2}\mathbf{g}_{2}^{H}&\cdots&\frac{c_{2}c_{M}}{S_{E,A}+1}\mathbf{g}_{2}\mathbf{g}_{M}^{H}\\ \cdots&\cdots&\cdots\\ \hline\cr\frac{c_{M}c_{2}}{S_{E,A}+1}\mathbf{g}_{M}\mathbf{g}_{2}^{H}&\cdots&\left(1+\frac{c_{M}^{2}}{S_{E,A}+1}\right)\mathbf{g}_{M}\mathbf{g}_{M}^{H}\end{array}\right]$ (167)
$=\left[\begin{array}[]{c|c}\mathbf{B}_{M-1}&\mathbf{C}_{M-1}\\ \hline\cr\mathbf{C}_{M-1}^{H}&\left(1+\frac{c_{M}^{2}}{S_{E,A}+1}\right)\mathbf{g}_{M}\mathbf{g}_{M}^{H}+\mathbf{I}\end{array}\right].$ (170)
Here $\mathbf{C}_{M-1}=\frac{c_{M}}{S_{E,A}+1}\mathbf{v}_{M-1}\mathbf{g}_{M}^{H}$.
We see that for $M=2$,
$t_{1,2}=\mathbf{v}_{2}^{H}\mathbf{B}_{2}^{-1}\mathbf{v}_{2}=\frac{c_{2}^{2}S_{E,2}}{(1+\frac{c_{2}^{2}}{S_{E,A}+1})S_{E,2}+1}$
$<\frac{c_{2}^{2}}{(1+\frac{c_{2}^{2}}{S_{E,A}+1})}<c_{2}^{2}.$ (171)
For $M>2$,
$\mathbf{B}_{M}^{-1}=\left[\begin{array}[]{cc}\mathbf{B}_{M|1,1}&\mathbf{B}_{M|1,2}\\ \mathbf{B}_{M|1,2}^{H}&\mathbf{B}_{M|2,2}\end{array}\right]$ (172)
where
$\mathbf{B}_{M|2,2}=\left(\left(1+\frac{c_{M}^{2}}{S_{E,A}+1}\right)\mathbf{g}_{M}\mathbf{g}_{M}^{H}+\mathbf{I}\right.$
$\,\,\left.-\mathbf{C}_{M-1}^{H}\mathbf{B}_{M-1}^{-1}\mathbf{C}_{M-1}\right)^{-1},$ (173)
$\mathbf{B}_{M|1,1}=\mathbf{B}_{M-1}^{-1}-\mathbf{B}_{M-1}^{-1}\mathbf{C}_{M-1}\mathbf{B}_{M|2,2}\mathbf{C}_{M-1}^{H}\mathbf{B}_{M-1}^{-1},$ (174)
$\mathbf{B}_{M|1,2}=-\mathbf{B}_{M-1}^{-1}\mathbf{C}_{M-1}\mathbf{B}_{M|2,2}.$ (175)
Then,
$t_{1,M}=v_{M}+2\Re\{w_{M}\}+\eta_{M}$ (176)
with $v_{M}=\mathbf{v}_{M-1}^{H}\mathbf{B}_{M|1,1}\mathbf{v}_{M-1}$, $w_{M}=\mathbf{v}_{M-1}^{H}\mathbf{B}_{M|1,2}c_{M}\mathbf{g}_{M}$, and $\eta_{M}=c_{M}^{2}\mathbf{g}_{M}^{H}\mathbf{B}_{M|2,2}\mathbf{g}_{M}$.
We know
$\mathbf{C}_{M-1}^{H}\mathbf{B}_{M-1}^{-1}\mathbf{C}_{M-1}=\frac{c_{M}^{2}}{(S_{E,A}+1)^{2}}\mathbf{g}_{M}\mathbf{v}_{M-1}^{H}\mathbf{B}_{M-1}^{-1}\mathbf{v}_{M-1}\mathbf{g}_{M}^{H}$
$=\frac{t_{1,M-1}c_{M}^{2}}{(S_{E,A}+1)^{2}}\mathbf{g}_{M}\mathbf{g}_{M}^{H},$ (177)
and hence
$\mathbf{B}_{M|2,2}=\left(\left(1+\frac{c_{M}^{2}}{S_{E,A}+1}-\frac{t_{1,M-1}c_{M}^{2}}{(S_{E,A}+1)^{2}}\right)\mathbf{g}_{M}\mathbf{g}_{M}^{H}+\mathbf{I}\right)^{-1}.$ (178)
Then
$\eta_{M}=\frac{c_{M}^{2}S_{E,M}}{\left(1+\frac{c_{M}^{2}}{S_{E,A}+1}-\frac{t_{1,M-1}c_{M}^{2}}{(S_{E,A}+1)^{2}}\right)S_{E,M}+1}$
$<\frac{c_{M}^{2}}{\left(1+\frac{c_{M}^{2}}{S_{E,A}+1}-\frac{t_{1,M-1}c_{M}^{2}}{(S_{E,A}+1)^{2}}\right)}.$ (179)
This bound is tight when $S_{E,M}$ is large. Also,
$v_{M}=\mathbf{v}_{M-1}^{H}\left(\mathbf{B}_{M-1}^{-1}\right.$
$\,\,\left.-\mathbf{B}_{M-1}^{-1}\mathbf{C}_{M-1}\mathbf{B}_{M|2,2}\mathbf{C}_{M-1}^{H}\mathbf{B}_{M-1}^{-1}\right)\mathbf{v}_{M-1}$
$=t_{1,M-1}-t_{1,M-1}^{2}\frac{c_{M}^{2}}{(S_{E,A}+1)^{2}}\mathbf{g}_{M}^{H}\mathbf{B}_{M|2,2}\mathbf{g}_{M}$
$=t_{1,M-1}-t_{1,M-1}^{2}\frac{\eta_{M}}{(S_{E,A}+1)^{2}}.$ (180)
$w_{M}=-\mathbf{v}_{M-1}^{H}\mathbf{B}_{M-1}^{-1}\mathbf{C}_{M-1}\mathbf{B}_{M|2,2}c_{M}\mathbf{g}_{M}$
$=-t_{1,M-1}\frac{c_{M}^{2}}{S_{E,A}+1}\mathbf{g}_{M}^{H}\mathbf{B}_{M|2,2}\mathbf{g}_{M}$
$=-t_{1,M-1}\frac{\eta_{M}}{S_{E,A}+1}.$ (181)
Therefore,
$t_{1,M}=t_{1,M-1}-t_{1,M-1}^{2}\frac{\eta_{M}}{(S_{E,A}+1)^{2}}$
$-2t_{1,M-1}\frac{\eta_{M}}{S_{E,A}+1}+\eta_{M}.$ (182)
We see $t_{1,M}<t_{1,M-1}+\eta_{M}<\sum_{i=2}^{M}\eta_{i}$.
It follows from (-C) that if $S_{E,A}+1>t_{1,i-1}$ for all $i$, we have $\eta_{i}<c_{i}^{2}$ for all $i$ (a tight bound when $S_{E,A}$ and $S_{E,i}$ are large), and hence
$t_{1,M}<\sum_{i=2}^{M}c_{i}^{2}<M-1.$ (183)
Here the first bound is tight when $S_{E,2},\cdots,S_{E,M}$ and $S_{E,A}$ are large, and the second bound is tight with additional large $S_{2},\cdots,S_{M}$. The above suggests that if $S_{E,A}$ is sufficiently large, then $t_{1,M}<M-1$. However, as shown next, $t_{1,M}$ is an increasing function of $S_{E,A}$, and hence $t_{1,M}<M-1$ for all $S_{E,A}$.
Note that it is easy to prove $\frac{\partial t_{1,M}}{\partial S_{E,A}}=-\mathbf{v}_{M}^{H}\mathbf{B}_{M}^{-1}\left(\frac{\partial}{\partial S_{E,A}}\mathbf{B}_{M}\right)\mathbf{B}_{M}^{-1}\mathbf{v}_{M}>0$ where $-\frac{\partial}{\partial S_{E,A}}\mathbf{B}_{M}$ is positive semi-definite. It is also easy to prove $\frac{\partial}{\partial S_{E,A}}\left(\frac{t_{1,M}}{S_{E,A}+1}\right)=-\mathbf{v}_{M}^{H}\mathbf{B}_{M}^{-1}\left(\frac{\partial}{\partial S_{E,A}}\left(\frac{1}{S_{E,A}+1}\mathbf{B}_{M}\right)\right)\mathbf{B}_{M}^{-1}\mathbf{v}_{M}<0$ where $\frac{\partial}{\partial S_{E,A}}\left(\frac{1}{S_{E,A}+1}\mathbf{B}_{M}\right)$ is positive semi-definite.
It follows from (-C) and (183) that
$\gamma_{1}-1=c_{1}-q_{1}$
$=\left(c_{1}-c_{1}^{2}\frac{S_{E,A}}{S_{E,A}+1}\right)-\frac{c_{1}^{2}t_{1,M}}{(S_{E,A}+1)^{2}}$ (184)
with $t_{1,M}<M-1$ for $M\geq 2$. This along with the initial discussion after (138) completes the proof of Proposition 5.